### v2

# smile: Single-turn to Multi-turn Inclusive Language Expansion via ChatGPT for Mental Health Support

Abstract
Developing specialized dialogue systems for mental health support requires multi-turn conversation data, which has recently garnered increasing attention. However, gathering and releasing large-scale and real-life multi-turn conversations to facilitate advancements in mental health presents challenges due to data privacy protection, as well as the time and cost involved. To address the challenges related to data scarcity, we introduce smile, a single-turn to multi-turn inclusive language expansion technique that prompts ChatGPT to rewrite public single-turn dialogues into multi-turn ones. Our work begins with the analysis of language transformation, validating the feasibility of the proposed method when compared with other baseline methods. We then conduct a study on dialogue diversity, including lexical features, semantic features, and dialogue topics, demonstrating the effectiveness of our proposed method. Furthermore, we implement an expert evaluation and the results demonstrate that the dialogues generated with our proposed method are of higher quality than those generated with other baseline methods. Thus, we employ our method to generate a large-scale, diverse, and high-quality dialogue dataset named SmileChat, comprising 55,165 dialogues in total with an average of 10.4 turns per dialogue. Finally, we utilize the collected corpus to develop a mental health chatbot, MeChat. To better assess the overall quality of SmileChat, we collect a real-life chat dataset comprising 82 counseling dialogues for model evaluation. Both automatic and human evaluations demonstrate that our trained dialogue system exhibits significant improvements, showcasing that SmileChat is high-quality and practical.
1 Introduction
Figure 1: Example of instructing ChatGPT to rewrite a single-turn dialogue into a multi-turn one.
We all know the importance of mental health and mental health issues (Kessler et al., 2005) have been a persistent concern for human beings. Recently, advancements in natural language processing (NLP) technology (Vaswani et al., 2017; Ouyang et al., 2022; Ni et al., 2022) have led to the emergence of neural-based conversational AI applied in various domains, including mental health (Liu et al., 2022; Tu et al., 2022). Virtual counselors powered by AI, as an innovative solution for mental health, can effectively address accessibility barriers, such as the high cost of treatment and the shortage of trained professionals to meet the demand. Furthermore, such a dialogue system provides mental health support as an effective and practical online counseling approach for those in need, safeguarding user privacy and mitigating the stigma that often accompanies help-seeking. However, the lack of publicly available, large-scale, diverse, and high-quality multi-turn chat datasets in the mental health support domain hinders the development of specialized dialogue systems.
Motivation Conversations related to mental health support often contain sensitive information and must be kept confidential (Lu et al., 2021) to safeguard the privacy of individuals seeking help. Making these conversations publicly available may discourage individuals from seeking support or negatively impact their personal and professional lives once known to people they are acquainted with. To facilitate progress in the NLP community, some researchers have attempted to collect various dialogue corpora (Liu et al., 2021; Sun et al., 2021; Zheng et al., 2022) through crowd-sourcing, data crawling, or data augmentation to build a dialogue agent capable of providing emotional and mental health support. How to construct a large-scale, diverse and high-quality multi-turn chat dataset for mental health motivates us to carry out the work as presented in this paper.
Challenges To be more specific, crowd-sourcing conversations (Liu et al., 2021) for emotional support has limitations due to the high cost and time required to train and manage annotators, as well as the difficulty in mimicking real-life interactions, that is, interlocutors may lack an understanding of the dilemma of living with mental disorders. An alternative is crawling QA (Sun et al., 2021) on a public mental health forum for training psychological support models. However, single-turn conversations may not be sufficient for resolving mental health issues, as multiple interaction exchanges are often needed. Multi-turn conversations, which can better simulate real-world conversations, are therefore more practical for training psychological support models. While the post-triggered machine-augmented method (Zheng et al., 2022) can address the limitations of scale and topic diversity, it does not take into account the responses of experienced supporters.
Our Approach To tackle the challenges mentioned above, we introduce smile, single-turn to multi-turn inclusive language expansion via ChatGPT. Specifically, we instruct ChatGPT to transform publicly available question-answer pairs (public QA), which can also be considered as single-turn dialogues from the real world, into multi-turn conversations, as illustrated in Figure 1. With the proposed method, we build a large-scale, diverse, and high-quality multi-turn conversation dataset for mental health support.
Our paper is organized as follows:
[…]
Our Contributions We make our data, code, and model publicly available. We believe our work offers a new perspective on constructing a large-scale, diverse, and high-quality multi-turn dialogue dataset for mental health within the research community. Our contributions can be summarized as follows:
  • We introduce smile, which provides a novel perspective for alleviating the scarcity of multi-turn conversations in mental health.
  • Through the analysis of language transformation and dialogue diversity, we verify the feasibility and effectiveness of our proposed method. This method can construct multi-turn dialogues based on medical, financial, and legal QAs, thereby alleviating the dialogue scarcity in other application domains.
  • To better assess the quality of SmileChat, we collect a real-life counseling dataset with 82 counseling dialogues to build an authentic test set, PsyTest, which contains 4045 utterances spoken by supporters. Both automatic and human evaluations demonstrate that SmileChat improves the performance of the dialogue system in mental health.
  • We release SmileChat, which comprises 55,165 Chinese multi-turn dialogues with an average of 10.4 turns. Additionally, we make our dialogue model, MeChat, and real-life test set, PsyTest, publicly available.
2 Related Work
2.1 Applications of ChatGPT
ChatGPT has proven to be a powerful AI tool for various NLP tasks since its release. Currently, it is being utilized in several domains, such as conversational AI (Alessa and Al-Khalifa, 2023; Köpf et al., 2023; Chen et al., 2023), education (Küchemann et al., 2023; Eshghie and Eshghie, 2023), code programming (Dong et al., 2023; Yetiştiren et al., 2023) and healthcare (Zhao et al., 2023; Yang et al., 2023).
Furthermore, ChatGPT’s efficiency and cost-effectiveness have been well-documented, making it competitive to human annotators (Gilardi et al., 2023; Zhu et al., 2023) even in zero-shot accuracy tasks. Xu et al. (2023) have proposed the use of self-chatting, where ChatGPT engages in a conversation with itself, resulting in 111.5k dialogues collected from Quora and Stack Overflow sources and 47k conversations from the medical domain. Auto-GPT  [^1] https://github.com/Significant-Gravitas/Auto-GPT, an AI agent, is capable of breaking down a natural language goal into sub-tasks and using various tools and the internet in an automated loop to achieve the objective. Shen et al. (2023) have suggested using ChatGPT for task planning when receiving user inquiries, selecting appropriate models based on function descriptions from Hugging Face, executing each subtask using the chosen AI model, and summarizing the response based on the execution’s outcomes.
In summary, ChatGPT has already demonstrated its enormous potential as an intelligent pipeline tool that can significantly advance NLP development, despite having only a restricted API available for researchers.
2.2 Datasets for Mental Health Support
Research on mental health support has significantly depended on the availability of publicly available datasets (Sun et al., 2021; Liu et al., 2021; Zheng et al., 2022) in recent years. The large-scale conversational datasets have enabled researchers to investigate various aspects of mental health, including identifying mental health conditions (Liu et al., 2023; Srivastava et al., 2022), understanding clients’ reactions (Li et al., 2023), predicting support strategies (Sun et al., 2021; Li et al., 2023), deciding personalized interventions (Golden et al., 2023) and understanding response safety within a dialogue history (Qiu et al., 2023).
Liu et al. (2021) first define the emotional support conversation task and then, via crowd-sourcing, construct ESConv, an emotional support conversation dataset containing 1053 dialogues with rich support strategies. However, the collection of ESConv requires high cost and time yet leads to a small-scale dialogue dataset. To this end, Zheng et al. (2022) present an approach for augmenting data scale with informative dialogue posts and then constructing AugESC, a model-synthesized dataset with 102k dialogues. The previous two datasets are limited to English. To facilitate the research in Chinese, hence Sun et al. (2021) crawl QA posts in a public mental health support platform compiling PsyQA.
3 Method
PsyQA [^2] https://www.xinli001.com/qa, a high-quality Chinese dialogue dataset related to mental health support in the form of one question mapping to multiple answers, has gone through a data anonymization process. Our dataset creation pipeline based on PsyQA comprises three main stages: (1) data preprocessing, (2) dialogue generation, and (3) dialogue retention.
3.1 Data Preprocessing
Considering the distinction between QAs in PsyQA and multi-turn dialogues and the context window limitation of 4096 tokens in ChatGPT, we recommend performing data preprocessing for PsyQA. This process involves wording cleaning and length truncation.
Wording Cleaning
This work aims to construct a large-scale, diverse, and high-quality multi-turn conversation corpus using the proposed smile method based on PsyQA. While QA can be considered a single-turn conversation between a real help-seeker and a supporter, there are differences in wording compared to actual multi-turn conversations. For instance, the term "楼主" (literally meaning "thread starter") frequently appears in QA but is rarely used in conversation. Therefore, we propose a two-stage process to clean the wording in PsyQA, mitigating linguistic discrepancies before rewriting QA into multi-turn conversations. This process includes both automatic and manual cleaning procedures. For detailed content, see Appendix A.
Length Truncation
After conducting a statistical analysis of the PsyQA dataset released by Sun et al. (2021), we find that 757 QAs had a total length exceeding 1800 characters. Additionally, we identify 9 QAs in which the total discourse length exceeds 4000 characters. Furthermore, the model gpt-3.5-turbo [^3] The model we use is gpt-3.5-turbo-0613. has a maximum context length of 4096 tokens. To ensure reliable and smooth rewrites, we limit the length of the QA pairs, maximizing the number of rewritten dialogue turns. Specifically, we truncate the length of the QA pairs at 1800 characters and truncate any excess text. This control ensures that the generated text is limited to approximately 2,000 tokens. It is worth noting that the PsyQA data used in this study undergo a data preprocessing process.
3.2 Dialogue Generation
3.2.1 Task Formulation
Considering the remarkable capabilities of ChatGPT and its cost-effective, plug-and-play characteristics, we consider ChatGPT to be a suitable choice for this task compared to other alternative models. Our study is designed to prompt ChatGPT to rewrite a single-turn dialogue into a multi-turn one. Generally, text generation can be formulated as follows:
$P_{\mathcal{M}_{\theta}}(Y|x)=\prod_{t=1}^{L}P_{\mathcal{M}_{\theta}}(y_{t}|Y_{<t},x)$ (1)
where $\mathcal{M}$ denotes a large language model with parameters $\theta$. $Y$ represents the output of the model generation, while $L$ signifies the variable length of the output. The symbol $y_{t}$ represents the $t$-th token generated by the model, and $x$ denotes the conditional model input.
3.2.2 Prompt Design
Figure 2: The smile method used to generate dialogues for mental health support.
To establish some conventions regarding specific terms for prompt design, we define the "help-seeker" and "supporter" as "求助者" and "支持者" in Chinese, respectively. A single-turn dialogue is defined as "求助者:$u^{H}$持者:$u^{S}$", where $u^{H}$ and $u^{S}$ represent the utterances of the help-seeker and supporter, respectively, and $H$ and $S$ refer to the help-seeker and supporter.
In this section, we mainly focus on detailing our methodology. Furthermore, we are building two baseline methods for comparison.
Standard Prompt
As its name suggests, the standard prompt does not contain any single-turn dialogues and instead uses only the simplest prompt to generate multi-turn dialogues. The specifics of the standard prompt are illustrated in Figure 8 in Appendix C. The conditional model input in Equation 1 is $x:=(\mathcal{I})$, where $\mathcal{I}$ represents the standard prompt. We simplify the method name as standard and consider this method as our baseline.
Standard Prompt with a Dialogue Topic
Intuitively, feeding a single, fixed prompt into a large language model often results in the generation of low diversity. To address this issue, we collaborate with three professional counselors, refer to existing literature (Rickwood et al., 2007; Pedrelli et al., 2015), and ultimately compile a comprehensive set of dialogue topics. This set comprises 56 distinct types, each accompanied by its corresponding explanation. For more details, please refer to Appendix K. The conditional model input in Equation 1 is $x:=(\mathcal{I,T})$, where $\mathcal{I}$ represents the dialogue topic chosen in uniform sampling. We simplify the method name to stand.${}_{\mathrm{DT}}$ and adopt this method as our baseline, as illustrated in Figure 8 in Appendix C.
smile Method
Our proposed method, referred to as the smile method, instructs the ChatGPT to rewrite single-turn dialogues into multi-turn ones. Figure 2 depicts the specific prompt template. The conditional model input in Equation 1 is $x:=(\mathcal{I,T,D})$, where $\mathcal{T}$ and $\mathcal{D}$ represent the dialogue topics hidden in the QA and single-turn dialogue, respectively.
3.3 Dialogue Retention
ChatGPT may exhibit potential instability during the dialogue generation process, but this occurs very infrequently. Therefore, we employ an automatic filtering mechanism to exclude dialogues that fail to meet our specified requirements, which encompass dialogue format and dialogue turns as detailed in Appendix B.
4 Language Transformation
4.1 Experimental Setup
QA Sampling
To ensure a fair comparison and prevent duplicate instances of the same question with different answers, we first randomly select 500 non-duplicate questions from the first 5,000 QAs in PsyQA. We then randomly choose one answer to serve as the corresponding question’s response. The data samples obtained are employed as seed dialogues, which are subsequently restructured into multi-turn conversations via ChatGPT. We name this 500 sampled data as PsyQA* to distinguish it from the whole dataset.
Hyperparameters
On the whole, we present three prompt methods in this paper. For each prompt method, we instruct ChatGPT to generate 500 dialogues to study language transformation and dialogue diversity, thereby validating the feasibility and effectiveness of our proposed smile method, respectively. To enhance the diversity of the generated dialogues, we set ChatGPT’s hyperparameters during text generation to the officially recommended default values, with $\tau$ (temperature) = 1.0 and $p$ (top_p) = 1.0.
Dialogue Filtering
[…]
Text Representation
Text representation is used for analyzing language transformation and semantic diversity. A dialogue between a help-seeker and a supporter is represented as $d=\{u_{0}^{H},u_{1}^{S},u_{2}^{H},...,u_{t}^{X},...,u_{n-1}^{H},u_{n}^{S}\}$, where $u_{t}^{X}$ represents the utterance of the $t$-th turn spoken by either the help-seeker or supporter. A string of a dialogue can be denoted as $d_{s}=[u_{0}^{H};u_{1}^{S};u_{2}^{H};...;u_{n-1}^{H};u_{n}^{S}]$, where [;][;][ ; ] denotes the operation of textual concatenation.
To obtain the text embedding of a dialogue, we employ OpenAI’s model text-embedding-ada-002444https://platform.openai.com/docs/guides/embeddings, which accepts a maximum context length of 8191. Each dialogue is first preprocessed into a single string without any speaker tokens and is then mapped to a 1536-dimensional vector. For example, to compute the cosine similarity between two different dialogues, we can obtain
$\mathrm{cos}(d_{i},d_{j})=\frac{E_{i}\cdot E_{j}}{\left\|E_{i}\right\|\left\|E_{j}\right\|}$ (2)
where $E_{i}$ and $E_{j}$ denote the text embeddings from two distinct dialogues.
4.2 Analysis
[…]
Figure 3 presents the distribution of dialogue transformation among three methods. According to our analysis, we conclude that single-turn dialogues can be successfully rewritten into multi-turn dialogues.
Figure 3: Distribution of dialogue transformation among three methods. The line $x=0.9312$ represents the boundary of $\mu-3\sigma$ in the smile method.
5 Dialogue Diversity
[…]
Figure 4: Pairwise dialogue cosine similarity among four settings: our proposed three methods and a reference point using sampled data from PsyQA.
5.1 Semantic Features
[…]
To calculate the pairwise cosine similarity between every two distinct dialogues, this produces $\binom{500}{2}$ pairs of dialogues and their corresponding cosine values in each prompt method, including the reference point of sampled data from PsyQA.
We present the results in Figure 4, which demonstrates that the median of the smile method is significantly lower than those of the baseline methods. The smile method exhibits the most extensive semantic diversity, aligning closely with the sampled dialogue of PsyQA. However, it’s worth noting that the median of the smile method is larger than that of PsyQA*. The reason behind this may be the introduction of token distribution from ChatGPT.
Category Total Help-seeker Supporter
# Dialogues 55165 - -
# Utterances 1833856 693756 1140100
# Avg. turns per dialogue 10.4 - -
Avg. utterances per dialogue 33.2 12.6 20.7
Avg. length per utterance 27.9 26.1 28.9
Table 1: Statistics of the dialogue dataset, SmileChat.
6 Expert Evaluation
We elaborate on data quality analysis from two main professional perspectives: (1) the evaluation of dialogue quality by three professional counselors, and (2) alignment with the supporters’ strategy flow patterns in real-life counseling conversations (For brevity, we include this section in Appendix G.).
To further validate that the dialogue quality generated using the SMILE method is superior to that produced by the two baseline methods, we propose to recruit three professional psychological counselors and randomly sample 20 dialogues for each method. Each set comprises three dialogues generated from different methods. Three counselors assign scores (3, 2, and 1, where 3 is the best, and 1 is the worst) based on three metrics: assessing (Q1) the overall quality of the conversation, (Q2) the authenticity of the user’s stated issues, and (Q3) the professionalism of the counselor’s responses.
We present the results of an expert evaluation of dialogue quality in 20 sets of dialogues in Table 2. Each value in the table represents the average score from three professional counselors, and results show that the dialogue quality generated with the SMILE method is of higher quality than that generated by the two baseline methods. The intraclass correlation coefficients (James et al., 1984; Koo and Li, 2016), including nine coefficients for the three metrics, are all greater than 0.8, indicating almost perfect agreement.
Method Q1 Q2 Q3
standard 1.95 2.12 1.21
stand.${}_{\mathrm{DT}}$ 1.25 1.2 1.82
smile 2.8 2.68 2.97
Table 2: Evaluation of dialogue quality by three professional counselors.
7 SmileChat Dataset
Through the analysis of language transformation, dialogue diversity and expert evaluation, we conclude that the proposed method can generate a diverse and high-quality chat dataset. Therefore, we utilize the smile method to guide ChatGPT in generating all multi-turn conversations based on PsyQA one round, leading to a large-scale dialogue dataset.
7.1 Data Statistics
To ensure data quality, we impose stricter requirements on dialogue turns, retaining only dialogues with at least 5 turns. Thus, we compile a collection of 55,165 conversations, SmileChat. Table 1 presents the statistics of the collected corpus.
7.2 Dialogue Exemplars
Multi-turn dialogue examples generated with the standard and stand.${}_{\mathrm{DT}}$ methods are illustrated in Figure 13 and 15. Further, a dialogue generated by the smile method is shown in Figure 17.
8 Dialogue System
We aim to build a high-quality multi-turn chat dataset for mental health. Therefore, we also analyze the dialogue quality based on the performance of the dialogue system trained with SmileChat.
8.1 Task Formulation
Our collected dataset can be represented as $D=\{d_{1},d_{2},...,d_{n}\}$, where each $d_{j}$ represents a single multi-turn dialogue.
To train a dialogue system for mental health, we need to split each dialogue into several training sessions. Specifically, a sampled $t$-turn dialogue session can be represented as follows:
$d_{t}=\left\{u^{H}_{0},u^{S}_{1},u^{H}_{2},...,u^{H}_{t-1},u^{S}_{t}\right\}\sim D$ (3)
We build a dialogue model that can predict the supporter’s utterance $u^{S}_{t}$ based on the dialogue history $h_{t}=\left\{u^{H}_{0},u^{S}_{1},u^{H}_{2},...,u^{H}_{t-1}\right\}$. Our objective is to maximize the likelihood probability as follows:
[…]
where $L$ is the sequence length of $u^{S}_{t}$.
8.2 Experimental Setup
Baseline Model
To validate the dialogue quality of our collected dataset, we conduct a fine-tuning experiment on ChatGLM2-6B (Zeng et al., 2023). Moreover, we also add two prompt engineering baselines (Prompt A and B, for more details, see Appendix I).
Parameter-efficient Fine-tuning
To preserve the original capabilities of the model while adapting to downstream dialogue tasks and reducing computational costs, we employ Low-Rank Adaptation (LoRA, (Hu et al., 2021)) on all linear layers in the ChatGLM2-6B model for efficient fine-tuning.
Hyperparameters
We present the hyperparameters for constructing a dialogue model for mental health in Table 10. For additional details on the training and generation processes, please refer to Appendix H.
Category Total Help-seeker Supporter
# Dialogues 82 - -
# Utterances 7672 3627 4045
# Avg. turns per dialogue 74.8 - -
Avg. utterances per dialogue 93.6 44.2 49.3
Avg. length per utterance 24.6 29.7 20.1
Table 3: Statistics of test set, PsyTest.
Methods PPL ($\Downarrow$) METEOR ($\Uparrow$) BLEU-1 ($\Uparrow$) BLEU-2 ($\Uparrow$) BLEU-3 ($\Uparrow$) Rouge-L ($\Uparrow$) D-1 ($\Uparrow$) D-2 ($\Uparrow$) D-3 ($\Uparrow$)
Baseline 3.38 ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$
Baseline + Prompt A 3.43 ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$
Baseline + Prompt B 3.99 ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$
Fine-tuned 1.50 ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$
Table 4: Results of automatic evaluation in our compiled real-life test set. $\dagger$ denotes the tokenizer we use is THUDM/chatglm2-6b, while $\ddagger$ denotes the tokenizer we use is hfl/chinese-roberta-wwm-ext-large.
8.3 Evaluation
To better understand and assess the dialogue quality of SmileChat dataset, we propose to utilize real-life multi-turn counseling conversations. We develop an online mental health support platform that enables professional counselors to offer each client a free text-based counseling service, lasting approximately 50 minutes each time. This duration is widely recognized as a standard time setting in psychological counseling. Therefore, we compile a collection of 82 high-quality real-life counseling dialogues with data desensitization for model evaluation. We name this test set PsyTest, which contains 4045 utterances spoken by supporters, as shown in Table 3.
8.3.1 Automatic Evaluation
Metrics
To conduct automatic evaluation, the evaluation metrics we use consist of Perplexity (PPL) (Jelinek et al., 1977), METEOR (Banerjee and Lavie, 2005), BLEU-1/2/3 (Papineni et al., 2002), Rouge-L (Lin, 2004), and Distinct-1/2/3 (D-1/2/3) (Li et al., 2016).
Results
The results of the automatic evaluation, including 9 metrics, are presented in Table 4. Notably, the evaluated dialogues are based on real-world counseling data rather than generated dialogues, which excludes the influence stemming from ChatGPT. All automatic evaluation metrics we use indicate improved performance. Our results show that the model trained with SmileChat is effective and practical. Consequently, the automatic evaluation demonstrates that our collected dataset is of high quality.
8.3.2 Human Evaluation
Metrics
We conduct a pairwise human evaluation to study the model performance trained with our proposed dialogue corpus. Initially, we randomly sample 100 multi-turn dialogue sessions, each comprising a minimum of 9 turns selected from PsyTest. Each dialogue session concludes with the utterance spoken by the counselor (also referred to as the supporter). Subsequently, we obtain 100 generated responses from the MeChat model. Three professional counselors are then presented with a dialogue history and three randomly shuffled responses (baseline, fine-tuned, ground truth). They are tasked with selecting the optimal response for the dialogue history, considering aspects such as naturalness, human-like qualities, correctness, and appropriateness. The evaluation is conducted based on the ethical principles of psychologists and the code of conduct (Association et al., 2016).
Figure 5: Pairwise human evaluation results, comparing the fine-tuned model to the baseline and ground truth, are reported. We present the win and lose rates of each compared pair in 100 randomly sampled real-life dialogue sessions. Fleiss’ kappa (Fleiss et al., 1981) is used to measure the inter-rater agreement, and all values fall within moderate agreement with $0.5\leq\kappa\leq 0.6$.
Results
We employ majority voting to reach final decisions among three professional counselors. As depicted in Figure 5, the model trained with SmileChat demonstrates a significant performance improvement compared to the baseline model. Moreover, the responses generated by MeChat surpass the golden response to some extent in 36% of all dialogue sessions.
Case Study
We present a case study, as shown in Figure 6. First, during counseling conversations, the text should not be too long for the help-seeker to understand. Second, at the beginning of a conversation, a supporter generally will not challenge the help-seeker; instead, they will provide support. Third, response 1 is more human-like and natural. Therefore, response 1 is better than response 2 overall.
Figure 6: Case study. Counseling strategies used in the two responses are presented in parentheses. Strategies in green are supporting, while those in red are challenging.
9 Conclusion
This paper introduces smile, a simple yet effective solution for addressing the scarcity of multi-turn conversations in mental health. Through language transformation and diversity analysis, we confirm the feasibility and effectiveness of our approach, demonstrating that our method is transferable. We also conduct expert evaluations, demonstrating that the smile method exhibits high superiority. Thus, the proposed method enables the automatic creation of a large-scale, diverse, and high-quality dialogue corpus, SmileChat, consisting of 55,165 dialogues with an average of 10.4 turns. Both automatic and human evaluations using the PsyTest dataset, consisting of 82 real-life counseling dialogues, demonstrate that SmileChat significantly improves dialogue system performance in mental health. With the release of multi-turn dialogues (SmileChat), our dialogue model (MeChat), and an authentic test set (PsyTest), we contribute valuable resources to the research community.
Limitations
We release a large-scale, diverse, and high-quality multi-turn conversational dataset for mental health support, generated by rewriting single-turn conversations into multi-turn conversations using ChatGPT. Consequently, the dataset unavoidably incorporates ChatGPT’s model knowledge. Furthermore, we discuss how ChatGPT does not fully utilize the rich vocabulary and content of single-turn conversations, as reflected in the distinct-$n$ metric.
Additionally, there is inevitably a gap in counseling practicality between models trained on this synthetic dataset and those trained on 50-minute counseling conversations, which will be our future work. While dialogues generated by the three methods still exhibit significant gaps when compared to the Xingling (Oracle) dataset with respect to the flow patterns of supporters’ strategies, there exists a large space for us to improve in our future work.
Furthermore, there is currently no comprehensive dataset available for evaluating the conversational safety of mental health support models. Therefore, we have identified this as a limitation of our paper and intend to address it in future work.
Ethical Considerations
We present a detailed illustration for the section on ethical considerations, which includes Data Usage Agreement, Usage Policy, and Suggestions for Real-life Deployment. Furthermore, the section on suggestions for real-life deployment highlights four main aspects: Anonymity and Consent, Handling Sensitive Topics, Ensuring Client Confidentiality, and Preventing Misinformation.
Data Usage Agreement
PsyQA is a single-turn dialogue dataset collected from an online mental health support platform. Specifically, help-seekers submit a post containing their mental health states and issues, and many professional counselors write down their responses to user questions to assist help-seekers. Therefore, PsyQA is a high-quality Chinese dialogue related to mental health support in the form of one question mapping to multiple answers.
Following the data copyright guidelines formulated by PsyQA (Sun et al., 2021), we release the multi-turn dialogue corpus publicly available for research purposes only. If researchers wish to reproduce the multi-turn dialogues using PsyQA, they must sign an agreement with the original data owner. Accordingly, we release our datasets and models for research purposes, thus facilitating further advancement in the academic community.
Usage Policy
Our chatbot is trained with the SmileChat dataset end-to-end. As illustrated in our paper, SmileChat still exhibits a significant gap in the flow patterns of counseling strategies compared to real-life counseling dialogues. Access to SmileChat and Mechat should be strictly limited to research purposes and should NOT be utilized for real-world deployment, commercial purposes, or any other use beyond academic research. Researchers using SmileChat and MeChat should be aware of the dataset and model limitations and make efforts to acknowledge and/or mitigate them to the best of their ability.
Suggestions for Real-life Deployment
When deploying our chatbot on the server for research purposes, we recommend implementing safeguards to protect help-seekers and, to a large extent, mitigate potential physical risks.
Anonymity and Consent
We strongly recommend that researchers allow users to interact with the chatbot anonymously, especially when discussing sensitive issues. Additionally, researchers should clearly communicate the level of anonymity and obtain informed consent from users before collecting any personal information.
Handling Sensitive Topics
In real-life human-machine interactions, we strongly recommend constructing a classifier to identify user intent. When users discuss sensitive topics, if detected by the intent classifier, a "human-in-the-loop" system allows a human moderator or healthcare professional to intervene in complex or sensitive situations.
Furthermore, we recommend ensuring clear escalation paths for users who require immediate human assistance, such as in cases of suicide and self-injury.
Ensuring Client Confidentiality
Researchers should clearly disclose to users that they are interacting with a chatbot, not a human, setting appropriate expectations for the system’s capabilities and limitations.
Researchers should provide clear information on how data is handled, stored, and for what purposes.
Misinformation Prevention
Researchers should implement fact-checking mechanisms to verify the accuracy of information provided by the chatbot.
Researchers should clearly communicate the limitations of the chatbot’s knowledge and advise users to consult a human professional for critical decisions.
Appendix A Details of Data Cleaning
A.1 Automatic Cleaning
We employ a sequential data cleaning pipeline to swiftly replace words that are unsuitable to the conversation scenario. For example, both "楼主你" (literally thread starter you) and "楼主" (literally thread starter) should be replaced with "你" (you). However, it is necessary to perform the former replacement to avoid the repetition of "你" and the resulting "你你" (you-you).
A.1.1 Word List for Data Cleaning
To avoid the repetition of "你" (you) and the resulting "你你" (you-you) , we suggest to conduct a sequential word replacing pipeline. Figure 7 shows the word list for data cleaning and corresponding order for automatic cleaning.
Figure 7: Word list for automatic cleaning.
A.2 Manual Cleaning
Due to the specificity and complexity of language, manual cleaning remains an essential part of the process. To prevent virtual dialogue systems from exhibiting overly frequent anthropomorphic behavior, we identify instances of the Chinese word for "hug" (抱抱) and manually delete sentence snippets containing this term.
Appendix B Requirements for Dialogue Retention
Here are two main requirements for dialogue filtering: data format and dialogue turns.
B.1 Data Format
We provide the requirements for data format as follows:
  • 1.
    The generated conversations do not start with "求助者:" or "支持者:".
  • 2.
[…]
  • 3.
    Each utterance in generated conversations does not start with "求助者:", "求助者:", "支持者:" or "支持者:".
  • 4.
    The last utterance in generated conversations contains an English sentence.
B.2 Dialogue Turns
Conversations comprising fewer than 5 turns will be discarded.
Appendix C Method
For brevity, we present the standard prompt and the standard prompt with a dialogue topic in Figure 8.
Figure 8: The standard prompt and the standard prompt with a dialogue topic. The two are separated by using the mark of "/" to differentiate between them. The key difference lies in whether a dialogue topic is introduced.
Appendix D Mechanism of Language Transformation
We propose two hypotheses: (1) When a single-turn dialogue is rewritten into a multi-turn dialogue, the similarity between the two is high (Attract). (2) When ChatGPT generates a multi-turn dialogue without introducing a single-turn dialogue, the similarity between the generated dialogue and a randomly selected single-turn dialogue is low (Repel). For better understanding, we present the mechanism of language transformation in Figure 9.
Figure 9: Mechanism for language transformation.
Appendix E Dialogue Topics Annotation
In this paper, to label the dialogue topics of generated dialogues, the hyperparameters of ChatGPT API during generation we used are set to the officially recommended default values, where temperature $\tau=1.0$ and $p=1.0$. Figure 10 shows the pipeline of dialogue topics annotation.
Figure 10: Pipeline of Dialogue Topics Annotation, where the content in bold is a placeholder.
Appendix F Dialogue Diversity
Metrics
# Unique
Unigrams
# Total
Unigrams
Distinct-1 ($\Uparrow$)
# Unique
Bigrams
# Total
Bigrams
Distinct-2 ()($\Uparrow$)( ⇑ )
# Unique
Trigrams
# Total
Trigrams
Distinct-3 ()($\Uparrow$)( ⇑ )
PsyQA* 11785 203306 0.058 120049 202806 0.592 182942 202306 0.904
standard 4174 153536 0.027 32281 153036 0.211 72340 152536 0.474
stand.${}_{\mathrm{DT}}$ 6032 175319 0.034 52141 174819 0.298 105921 174319 0.608
smile 10447 254585 0.041 111662 254085 0.439 196367 253585 0.774
Table 5: Statistics of 500 conversations in each prompt method, including PsyQA*.
F.1 Lexical Features
For lexical analysis, we utilize the ChatGLM2-6B tokenizer [^5] https://huggingface.co/THUDM/chatglm2-6b to tokenize the dialogue. To measure the lexical features, we adopt distinct-$n$ ($n=1,2,3$) metrics (Li et al., 2016), which are widely used for measuring the diversity of dialogue datasets. Each dialogue is first preprocessed into a single string without any speaker tokens. We provide statistics for 500 dialogues per prompt method, with PsyQA* serving as the reference point, detailed in Table 5.
Our proposed smile method results in rich vocabularies, with a significantly higher number of unique unigrams, bigrams and trigrams compared to the baseline methods. Specifically, a simple and fixed prompt tends to produce monotonous output. When dialogue topics are incorporated into a single, fixed prompt, the model’s output demonstrates substantial diversification. Furthermore, the smile method outperforms the baseline methods in terms of Distinct-1, Distinct-2 and Distinct-3.
F.2 Dialogue Topics
To measure the diversity of dialogue topics in a dialogue dataset, we utilize information entropy to measure the diversity of topic distribution. The higher the information entropy, the more uniform the distribution, indicating greater diversity. The formula for calculating information entropy (Rényi, 1961; Lin, 1991) is as follows:
$H(X)=-\sum_{i=1}^{n}p(x_{i})\log_{2}{p(x_{i})}$ (5)
where $H(X)$ is the information entropy. $p(x_{i})$ is the probability of the occurrence of topic $x_{i}$.
To obtain dialogue topics for each dialogue in each prompt method, we design a prompt provided with 56 distinct dialogue topics, as illustrated in Appendix K and Figure 10, to automatically label dialogue topics for each dialogue with GPT-4. The model we use is gpt-4-0613. We present the information entropy for each prompt method in Table 6, demonstrating that the dialogues generated using the smile method are substantially more diverse than those generated using the standard method and are compatible with the stand.${}_{\mathrm{DT}}$ method, which uniformly samples dialogue topics.
Setting standard stand.${}_{\mathrm{DT}}$ smile
Limited Topics 8.11 14.28 14.07
Unlimited Topics 8.40 14.76 15.02
Average 8.26 14.52 14.55
Table 6: Information entropy of dialogue topics. The setting of limited topics indicates that the dialogue topics in the statistics are confined to the 56 dialogue topics we provide. The setting of unlimited topics denotes that the dialogue topics not only belong to our provided 56 distinct topics but also include topics generated spontaneously by GPT-4.
Appendix G Expert Evaluation
To illustrate the patterns of supporters’ strategies, we compile statistics on the flow patterns of these strategies, including "Supporting", "Challenging" and "Other".
G.1 Experimental Setup
PsyQA (Sun et al., 2021) contains 4012 QAs with supporters’ strategies in the original dataset. Xinling (Li et al., 2023) is a real-life multi-turn counseling dataset containing 550 dialogues with supporters’ strategies in the original dataset.
Three methods (standard, stand.${}_{\mathrm{DT}}$, and smile) each contain 500 dialogues, respectively. We use Xinling to train a classifier to predict the supporters’ strategies in dialogues generated with these three methods. The performance of the trained classifier is comparable to the results in the original paper.
G.2 Results
Supporters have three main strategies: "Supporting", "Challenging", and "Other", and we present the results in Table 7, 8 and 9. We highlight the gap greater the 0.2 in bold and the corresponding gap is in parentheses.
Strategy Supporting
Methods standard stand.DT SMILE PsyQA Xinling (Oracle)
Stage 1 0.854(+0.101) 0.847(+0.094) 0.615(-0.138) 0.474(-0.279) 0.753
Stage 2 0.547(-0.278) 0.520(-0.305) 0.432(-0.393) 0.312(-0.513) 0.825
Stage 3 0.419(-0.367) 0.365(-0.421) 0.360(-0.426) 0.179(-0.607) 0.786
Stage 4 0.468(-0.258) 0.387(-0.339) 0.390(-0.336) 0.107(-0.619) 0.726
Stage 5 0.808(+0.382) 0.774(+0.348) 0.662(+0.236) 0.154(-0.272) 0.426
Table 7: Pattern flow of "Supporting" strategy.
Strategy Challenging
Methods standard stand.DT SMILE PsyQA Xinling(Oracle)
Stage 1 0.140(+0.084) 0.150(+0.094) 0.372(+0.316) 0.348(+0.292) 0.056
Stage 2 0.452(+0.305) 0.478(+0.331) 0.565(+0.418) 0.598(+0.451) 0.147
Stage 3 0.576(+0.384) 0.632(+0.440) 0.636(+0.444) 0.752(+0.560) 0.192
Stage 4 0.516(+0.272) 0.603(+0.359) 0.586(+0.342) 0.847(+0.603) 0.244
Stage 5 0.112(-0.127) 0.157(-0.082) 0.231(-0.008) 0.794(+0.555) 0.239
Table 8: Pattern flow of "Challenging" strategy.
Strategy Other
Methods standard stand.DT SMILE PsyQA Xinling(Oracle)
Stage 1 0.006(-0.185) 0.003(-0.188) 0.013(-0.178) 0.178(-0.013) 0.191
Stage 2 0.001(-0.027) 0.002(-0.026) 0.003(-0.025) 0.090(+0.062) 0.028
Stage 3 0.005(-0.017) 0.003(-0.019) 0.004(-0.018) 0.069(+0.047) 0.022
Stage 4 0.016(-0.014) 0.010(-0.020) 0.024(-0.006) 0.046(+0.016) 0.03
Stage 5 0.080(-0.255) 0.069(-0.266) 0.107(-0.228) 0.052(-0.283) 0.335
Table 9: Pattern flow of "Other" strategy.
G.3 Conclusion
Dialogues generated by the three methods exhibit similar flow patterns of supporters’ strategies. Instructing ChatGPT to rewrite QAs into multi-turn dialogues will inevitably impact the strategy pattern in the generated dialogues. Dialogues generated by the three methods still exhibit significant gaps when compared to the Xingling (Oracle) dataset with respect to the flow patterns of supporters’ strategies.
Appendix H Implementation Details
Model Training.
We implement ChatGLM2-6B with the Transformers library. Table 10 shows the parameters of parameter-efficient fine-tuning.
Epoch
Learning
Rate
Batch
Size
LoRA
Rank
LoRA
Dropout
LoRA
$\alpha$
Seed
2 1e-4 1 16 0.1 64 1234
Table 10: Parameters of parameter-efficient fine-tuning.
Dialogue Generation
We use ChatGLM2-6B as the baseline model, and all hyperparameters during generation are set to their default values from the official repository [^6] https://huggingface.co/THUDM/chatglm2-6b. After tuning ChatGLM2-6B with LoRA, we set the maximum generation length to 2040, temperature to 0.8, and adopt nucleus sampling with $p=0.8$.
Appendix I Prompt Details
Prompt A is directly utilized from an online GitHub repository, while prompt B is designed by our professional counselors.
Prompt A [^7] https://github.com/PlexPt/awesome-chatgpt-prompts-zh
我想让你担任心理医生。我将为您提供一个寻求指导和建议的人,以管理他们的情绪、压力、焦虑和其他心理健康问题。您应该利用您的认知行为疗法、冥想技巧、正念练习和其他治疗方法的知识来制定个人可以实施的策略,以改善他们的整体健康状况。 (English translation: I want you to take on the role of a psychologist. I will provide you with someone seeking guidance and advice to manage their emotions, stress, anxiety, and other mental health issues. You should use your knowledge of cognitive-behavioral therapy, meditation techniques, mindfulness practices, and other therapeutic methods to formulate strategies that individuals can implement to improve their overall health.)
Prompt B
现在你扮演一位专业的心理咨询师,你具备丰富的心理学和心理健康知识。你擅长运用多种心理咨询技巧,例如认知行为疗法原则、动机访谈技巧和解决问题导向的短期疗法。以温暖亲切的语气,展现出共情和对来访者感受的深刻理解。以自然的方式与来访者进行对话,避免过长或过短的回应,确保回应流畅且类似人类的对话。提供深层次的指导和洞察,使用具体的心理概念和例子帮助来访者更深入地探索思想和感受。避免教导式的回应,更注重共情和尊重来访者的感受。根据来访者的反馈调整回应,确保回应贴合来访者的情境和需求。请为以下的对话生成一个回复。 (English translation: Now, you are playing the role of a professional psychological counselor with extensive knowledge of psychology and mental health. You excel in applying various counseling techniques, such as principles of cognitive-behavioral therapy, motivational interviewing skills, and solution-focused short-term therapy. With a warm and friendly tone, demonstrate empathy and a profound understanding of the visitor’s feelings. Engage in a natural conversation with the visitor, avoiding overly long or short responses to ensure a smooth and human-like dialogue. Provide deep guidance and insights, using specific psychological concepts and examples to help the visitor explore thoughts and feelings more deeply. Avoid instructive responses and focus more on empathy and respecting the visitor’s feelings. Adjust responses based on visitor feedback to ensure they align with the visitor’s context and needs. Please generate a response for the following dialogue.)
Appendix J Instructions for Human Evaluation
The three professional counselors are willing to help and are interested in this research. Furthermore, their average age is 30 years old, with two females and one male among them. We present our instructions for human evaluation in Figure 11.
Figure 11: Labeling instruction.
To maintain the fairness of model evaluation, three responses randomly appear in a different order every time. Furthermore, three professional psychologists are willing to evaluate the response quality, ensuring the quality of human evaluation.
Pairwise human evaluation results, comparing the baseline model to ground truth, are reported, as illustrated in Figure 12. Results show that the baseline model, without training with SmileChat, lags significantly behind compared to the ground truth.
Figure 12: Pairwise human evaluation results, comparing the baseline model to ground truth, are reported. We present the win and lose rates of each compared pair in 100 randomly sampled real-life dialogue sessions. Fleiss’ kappa (Fleiss et al., 1981) is used to measure the inter-rater agreement, and the value falls within moderate agreement with $0.5\leq\kappa\leq 0.6$.
Appendix K Definitions of Dialogue Topics
1. Mate Selection: Do not know how to enter into an intimate relationship, have fear or aversion towards intimate relationships, are unwilling to date, or encounter some difficulties in choosing a spouse.
2. Love Issues: Problems encountered by individuals in a romantic state during the process of love, including both long-distance and non-long-distance relationships.
3. Post-love Issues: Interpersonal relationship management with ex-partners after the end of an intimate relationship.
4. Marriage Issues: Limit to relationship problems between spouses. Marriage issues refer to a series of problems that arise between a husband and wife after marriage. They generally include extramarital affairs, emotional infidelity, domestic violence, personality defects, personality disorders, mother-in-law and daughter-in-law relationships, dull marriage, interpersonal communication, cultural differences, sexual life, and problems in marriages caused by practices like polygamy or polyandry.
5. Concepts of Sex Distress: Sexual concept distress refers to the distress present in understanding and views regarding sexual physiology, psychology, behavior, morality, and civilization.
6. Sexual Preference Distress: Sexual preference distress refers to the psychological distress caused by paraphilias such as fetishism, transvestic fetishism, exhibitionism, friction fetishism, voyeurism, zoophilia, pedophilia, sadomasochism, and necrophilia.
7. Gender Identity/Cognition: It refers to the deeply felt gender within a person, based on personal experience, which may align with the gender assigned at birth (i.e., cisgender) or differ from it (i.e., transgender).
8. Sexual Orientation: Refers to the gender(s) a person is attracted to (including the opposite sex, same-sex, no specific sex, both sexes).
9. Family Conflict: Conflicts, disputes, and communication issues among family members, including those originating from both the nuclear family and the extended family.
10. Child Education: Literally meaning the education of children.
11. Domestic Violence: Physically, mentally, or otherwise harmful actions carried out between family members, including acts such as physical assault, binding, harming, and restricting personal freedom, as well as consistent verbal abuse, intimidation, and similar behaviors.
12. Sexual Harassment: Sexual harassment refers to using sexually suggestive language or actions towards the target of harassment, coercing the victim to comply and making them uncomfortable.
13. Sexual Assault: Sexual assault involves various forms of unwanted sexual contact and coerced sexual activities, including rape, forced kissing, sexual harassment, sexual abuse, exhibitionism, voyeurism, and such actions might also be considered sexual assault in legal precedents.
14. Bullying: Typically refers to the harassment and oppression between individuals with unequal power, which has long existed in society. This includes physical or verbal attacks, resistance and exclusion in interpersonal interactions. It can also involve talking about someone in a sexual harassment-like manner or mocking and making comments about their body parts, or it could be insults and sarcasm driven by personal reasons like jealousy.
15. Loss: The loss of significant others or pets.
16. Setbacks: Inevitable emotional reactions resulting from the obstruction of purposeful actions by individuals. It can cause substantial harm and manifest as disappointment, pain, distress, and unease.
17. Political Violence: Rooted in class domination and relying on the state as its entity, political coercion is the force exerted by the state, relying on various means of coercion such as the military, police, courts, and prisons.
18. Secondary Trauma: After witnessing a large number of cruel and destructive scenes, the degree of psychological and emotional distress exceeds the tolerance limit of some individuals, leading to various abnormal psychological phenomena. These phenomena are often a result of sympathy and empathy for survivors and their trauma, causing severe mental and emotional distress and even psychological breakdown. The main symptoms of secondary trauma include loss of appetite, easy fatigue, decreased physical energy, sleep disturbances (difficulty falling asleep, easy awakening), nightmares, irritability or anger, easy startle response, lack of concentration, feeling numb, fearful, and hopeless about one’s own experiences, accompanied by trauma reactions and interpersonal conflicts.
19. Trauma from Major Life Events: Psychological shadows caused by significant life events other than loss.
20. Psychological Counseling Trauma: Harm caused by therapists during psychological counseling.
21. Health Issues: Life and psychological distress caused by diseases such as heart disease, thyroid nodules, polycystic ovary syndrome, et al.
22. Psychosomatic Symptoms: Including palpitations (a subjective feeling of discomfort with heartbeats), sleep problems (such as insomnia/nightmares), which, if caused by sleep issues leading to other problems, are categorized as sleep problems; eating issues (such as loss of appetite/emotional eating); memory problems; stomach pain; dizziness; fainting; difficulty breathing; physical weakness/tiredness without apparent reason; unexplained bodily pains et al.
23. School/Workplace Adaptation: Literal meaning.
24. Role Transition Adaptation: For example, new mothers/fathers, new wives/husbands, retirees.
25. Cultural Adaptation: Adaptive issues encountered during the ongoing direct contact between individuals from different cultural groups in the process of changes occurring in one or both parties due to their original cultural backgrounds.
26. Self-exploration and Growth: Self-exploration and growth in the course of life development (adolescence, early adulthood, middle age, late adulthood).
27. Personality Trait Exploration: Traits, formation, origins, influences, et al., of personality and character.
28. Exploration of Negative Self-evaluation: Not knowing how to love oneself, feelings of inferiority, low self-esteem, self-denial, self-doubt, internal conflicts, self-contradictions, fearing being different, sensitivity, lack of security, feeling inadequate, et al.
29. Exploration of Life Meaning: Sense of meaninglessness, existential emptiness.
30. Emotion Regulation/Control Methods Consultation: Literal meaning.
31. Depression: An emotion characterized by "low mood, slow thinking, reduced speech and movement."
32. Anxiety: A restless emotion arising from excessive worry about the safety of loved ones, one’s own life, future, destiny, et al.
33. Stress: Stress is a cognitive and behavioral experiential process formed by stressors and stress responses, namely psychological stress.
34. Obsession: Marked by an excessive pursuit of perfection, precision, and a tendency to rationalize conflicts. It involves strong self-control tendencies and self-regulatory behavior to the point of entanglement and nitpicking. Behaviorally, it adheres excessively to rules, regulations, and orders, to the extent that even life details are sought to be procedural and ritualistic, demanding step-by-step compliance.
35. Fear: A strong and repressed emotional state felt deeply by an individual or group in the face of real or imagined danger. Manifestations include heightened nervousness, overwhelming inner fear, inability to concentrate, mental blankness, impaired judgment or self-control, and increased impulsiveness.
36. Dilemma/Decision-Making Difficulty: Faced with numerous choices and concerns, dealing with dilemmas or even multiple predicaments, struggling to make decisions.
37. Impulsiveness/Loss of Control: Often refers to acting recklessly without considering consequences. It involves intense emotions and weak rational control.
38. Interpersonal Communication Skills/Methods Consultation: Inquire about handling interpersonal relationships and more.
39. Interpersonal Conflicts/Disputes: Conflicts, disagreements, dissatisfaction, or communication issues arise during interpersonal interactions. If there are other issues caused by interpersonal relationships, they are also categorized as interpersonal interactions.
40. Social Difficulties/Fears: Characterized by involuntarily feeling nervous and frightened when interacting with others (especially in public settings), leading to confusion, incoherent speech, and even severe fear of social encounters.
41. Learning Efficiency/Methods: Literal meaning.
42. Work Efficiency/Methods: Literal meaning.
43. Job Dissatisfaction: Includes dissatisfaction with salary, benefits, environment, systems, personnel, et al. in a work setting.
44. Learning Dissatisfaction: Includes dissatisfaction with interpersonal relationships, environment, systems, et al. in an educational setting.
45. Occupational Burnout: Refers to the state of physical and mental fatigue and exhaustion experienced by individuals under heavy work pressure.
46. Learning Fatigue: Refers to a phenomenon where students hold a negative attitude towards school courses and academics, accompanied by the following behavioral manifestations: loss of enthusiasm for academic tasks and school activities, displaying a negative state, and showing indifference and estrangement towards classmates and friends.
47. Occupational Stress/Challenges: Various pressures/challenges that arise or form in the workplace, including stress/challenges caused by factors such as heavy workload, difficulties in interpersonal communication, and the impact of changes in the work environment.
48. Academic Pressure: Refers to the mental burden that individuals bear during learning activities. This includes various tense stimuli from the environment experienced during the learning process, as well as measurable and assessable abnormal reactions in the physiological, psychological, and social behavior of students.
49. Dropping Out: Stopping attending school midway or losing the opportunity to attend school.
50. Underemployment/Unemployment: Underemployment refers to the behavior of not finding a job and waiting for job opportunities. Unemployment refers to the situation of a person within a certain age range who is willing and capable of working for remuneration but has yet to find a job.
51. Career Planning: Refers to the ongoing systematic process of planning for one’s career and even life. It includes three elements: career orientation, goal setting, and pathway design.
52. Professional/Career/Occupational Choices/Changes: Literal meaning.
53. Suspected Neuroses and Mental Disorders: Highly suspected of having neuroses and mental disorders, recommended type of consultation.
54. Neuroses and Mental Disorders: It is a general term for a group of mental disorders, including neurasthenia, obsessive-compulsive disorder, anxiety disorder, phobia, somatic symptom disorder, et al. Patients experience significant distress and impairment in psychological or social functioning, but there is no confirmed organic pathological basis. The course of the illness is mostly prolonged or episodic. Mid- to long-term consultations are needed. Mental disorders refer to a range of conditions that meet the diagnostic criteria of DSM-5 (substance addiction or dependence also included) and require hospital consultation.
55. Self-harm/Suicidal Tendencies: Describing incidents of suicide/self-harm/self-mutilation/contemplating suicide. Triggers crisis alerts and referrals.
56. Harming Others/Killing: Describing incidents of harming others/killing/thinking of killing.
Appendix L Dialogue Examplars
Multi-turn dialogue examples generated with the standard and stand.${}_{\mathrm{DT}}$ methods are illustrated in Figure 13 and 15. Further, a dialogue generated by the smile method is shown in Figure 17.
Figure 13: A dialogue example generated with the standard method. (Chinese Version)
Figure 14: A dialogue example generated with the standard method. (English Version)
Figure 15: A dialogue example generated with the stand.${}_{\mathrm{DT}}$ method. (Chinese Version)
Figure 16: A dialogue example generated with the stand.${}_{\mathrm{DT}}$ method. (English Version)
Figure 17: An example of multi-turn dialogue generated by smile method. (Chinese Version)
Figure 18: An example of multi-turn dialogue generated by smile method. (English Version)

### v3

@@
-smile: Single-turn to Multi-turn Inclusive Language Expansion via ChatGPT for Mental Health Support
+Huachuan Qiu1,2 Hongliang He1,2 Shuai Zhang1,2 Anqi Li1,2 Zhenzhong Lan2,3
 Abstract
 Developing specialized dialogue systems for mental health support requires multi-turn conversation data, which has recently garnered increasing attention.
-However, gathering and releasing large-scale and real-life multi-turn conversations to facilitate advancements in mental health presents challenges due to data privacy protection, as well as the time and cost involved.
-To address the challenges related to data scarcity, we introduce smile, a single-turn to multi-turn inclusive language expansion technique that prompts ChatGPT to rewrite public single-turn dialogues into multi-turn ones.
-Our work begins with the analysis of language transformation, validating the feasibility of the proposed method when compared with other baseline methods.
-We then conduct a study on dialogue diversity, including lexical features, semantic features, and dialogue topics, demonstrating the effectiveness of our proposed method.
-Furthermore, we implement an expert evaluation and the results demonstrate that the dialogues generated with our proposed method are of higher quality than those generated with other baseline methods.
-Thus, we employ our method to generate a large-scale, diverse, and high-quality dialogue dataset named SmileChat, comprising 55,165 dialogues in total with an average of 10.4 turns per dialogue.
+However, gathering and releasing large-scale, real-life multi-turn conversations that could facilitate advancements in mental health support presents challenges in data privacy protection and the time and cost involved in crowdsourcing.
+To address these challenges, we introduce SMILE, a single-turn to multi-turn inclusive language expansion technique that prompts ChatGPT to rewrite public single-turn dialogues into multi-turn ones.
+Our work begins by analyzing language transformation and validating the feasibility of our proposed method.
+We conduct a study on dialogue diversity, including lexical features, semantic features, and dialogue topics, demonstrating the effectiveness of our method.
+Further, we employ our method to generate a large-scale, lifelike, and diverse dialogue dataset named SMILECHAT, consisting of 55k dialogues.
 Finally, we utilize the collected corpus to develop a mental health chatbot, MeChat.
-To better assess the overall quality of SmileChat, we collect a real-life chat dataset comprising 82 counseling dialogues for model evaluation.
-Both automatic and human evaluations demonstrate that our trained dialogue system exhibits significant improvements, showcasing that SmileChat is high-quality and practical.
+To better assess the quality of SMILECHAT, we collect a small-scale real-life counseling dataset conducted by data anonymization.
+Both automatic and human evaluations demonstrate significant improvements in our dialogue system and confirm that SMILECHAT is high-quality.
+Code, data, and model are publicly available at https://github.com/qiuhuachuan/smile.
 1 Introduction
-Figure 1: Example of instructing ChatGPT to rewrite a single-turn dialogue into a multi-turn one.
-We all know the importance of mental health and mental health issues (Kessler et al., 2005) have been a persistent concern for human beings.
-Recently, advancements in natural language processing (NLP) technology (Vaswani et al., 2017; Ouyang et al., 2022; Ni et al., 2022) have led to the emergence of neural-based conversational AI applied in various domains, including mental health (Liu et al., 2022; Tu et al., 2022).
-Virtual counselors powered by AI, as an innovative solution for mental health, can effectively address accessibility barriers, such as the high cost of treatment and the shortage of trained professionals to meet the demand.
-Furthermore, such a dialogue system provides mental health support as an effective and practical online counseling approach for those in need, safeguarding user privacy and mitigating the stigma that often accompanies help-seeking.
+We all know the importance of mental health, and mental health issues have been a persistent concern for human beings (Kessler et al., 2005).
+Recent advancements in natural language processing (NLP) technology (Ouyang et al., 2022; Ni et al., 2022) have led to the emergence of neural-based conversational AI in mental health support (Liu et al., 2022; Tu et al., 2022).
+As an innovative solution for mental health, virtual counselors powered by large language models (LLMs) can effectively address accessibility barriers, such as the high cost of treatment and the shortage of experienced professionals to meet the demand.
+Furthermore, such dialogue systems provide mental health support as an effective and practical online counseling approach for those in need, safeguarding user privacy and mitigating the stigma that often accompanies help-seeking.
 However, the lack of publicly available, large-scale, diverse, and high-quality multi-turn chat datasets in the mental health support domain hinders the development of specialized dialogue systems.
-Motivation Conversations related to mental health support often contain sensitive information and must be kept confidential (Lu et al., 2021) to safeguard the privacy of individuals seeking help.
-Making these conversations publicly available may discourage individuals from seeking support or negatively impact their personal and professional lives once known to people they are acquainted with.
-To facilitate progress in the NLP community, some researchers have attempted to collect various dialogue corpora (Liu et al., 2021; Sun et al., 2021; Zheng et al., 2022) through crowd-sourcing, data crawling, or data augmentation to build a dialogue agent capable of providing emotional and mental health support.
-How to construct a large-scale, diverse and high-quality multi-turn chat dataset for mental health motivates us to carry out the work as presented in this paper.
-Challenges To be more specific, crowd-sourcing conversations (Liu et al., 2021) for emotional support has limitations due to the high cost and time required to train and manage annotators, as well as the difficulty in mimicking real-life interactions, that is, interlocutors may lack an understanding of the dilemma of living with mental disorders.
+Figure 1: The SMILE method used to generate dialogues for mental health support.
+Motivation
+Indeed, many researchers have been pursuing the goal of building a practical and effective conversational agent for mental health.
+However, the first step in creating such a system is to have training data.
+Conversations related to mental health support often contain sensitive information and must be kept confidential to safeguard the privacy of individuals seeking help (Lu et al., 2021).
+Making these conversations publicly available may discourage individuals from seeking support or negatively impact their personal and professional lives once known to people with whom they are acquainted.
+To facilitate progress in the NLP community, some researchers have attempted to collect various dialogue corpora (Liu et al., 2021; Sun et al., 2021; Zheng et al., 2022) through crowdsourcing, data crawling, or data augmentation to build a dialogue agent capable of providing emotional and mental health support.
+How to construct a large-scale, diverse, and high-quality multi-turn chat dataset for mental health without effort motivates us to carry out the work as presented in this paper.
+Challenges
+To be more specific, crowdsourcing conversations (Liu et al., 2021) for emotional support has limitations due to the high cost and time required to train and manage annotators, as well as the difficulty in mimicking real-life interactions, that is, interlocutors may lack an understanding of the dilemma of living with mental disorders.
 An alternative is crawling QA (Sun et al., 2021) on a public mental health forum for training psychological support models.
 However, single-turn conversations may not be sufficient for resolving mental health issues, as multiple interaction exchanges are often needed.
-Multi-turn conversations, which can better simulate real-world conversations, are therefore more practical for training psychological support models.
-While the post-triggered machine-augmented method (Zheng et al., 2022) can address the limitations of scale and topic diversity, it does not take into account the responses of experienced supporters.
-Our Approach To tackle the challenges mentioned above, we introduce smile, single-turn to multi-turn inclusive language expansion via ChatGPT.
-Specifically, we instruct ChatGPT to transform publicly available question-answer pairs (public QA), which can also be considered as single-turn dialogues from the real world, into multi-turn conversations, as illustrated in Figure 1.
+Multi-turn conversations, which can better simulate real-world conversations, are more practical for training psychological support models.
+While the post-triggered machine-augmented method (Zheng et al., 2022) can address the limitations of scale and topic diversity, it does not consider the responses of experienced supporters.
+Our Approach
+To tackle the abovementioned challenges, we introduce SMILE, single-turn to multi-turn inclusive language expansion via ChatGPT, as shown in Figure 1.
+Specifically, we instruct ChatGPT to transform publicly available long question-answer pairs (public QAs), which can also be viewed as single-turn dialogues, into multi-turn conversations.
 With the proposed method, we build a large-scale, diverse, and high-quality multi-turn conversation dataset for mental health support.
 Our paper is organized as follows:
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Our Contributions We make our data, code, and model publicly available.
-We believe our work offers a new perspective on constructing a large-scale, diverse, and high-quality multi-turn dialogue dataset for mental health within the research community.
+  • We first present our method (§$\S$3), including data preparation, task definition, and prompt design that elaborates on the SMILE method and other baseline methods.
+  • We then demonstrate the feasibility of the SMILE method through language transformation (§$\S$4), showing that the dialogue constructed by this method is lifelike.
+  • Subsequently, we demonstrate the effectiveness of the SMILE method by utilizing three diversity indicators (§$\S$5): lexical features, semantic features, and dialogue topics.
+Following the validation of feasibility and effectiveness, we leverage the SMILE method to generate a large-scale, lifelike, and diverse multi-turn chat dataset, SMILECHAT.
+  • Finally, we propose to train a dialogue system to explore the quality of conversation (§$\S$6) and collect a set of 50 anonymized real-life counseling dialogues for model evaluation.
+Our Contributions
+We make our data, code, and model publicly available.
+We believe that our work offers a new perspective on constructing a large-scale, lifelike, diverse, and high-quality multi-turn dialogue dataset for mental health within the research community.
 Our contributions can be summarized as follows:
-  • We introduce smile, which provides a novel perspective for alleviating the scarcity of multi-turn conversations in mental health.
+  • We introduce SMILE, which provides an easy and novel method for alleviating the scarcity of multi-turn conversations in mental health support.
   • Through the analysis of language transformation and dialogue diversity, we verify the feasibility and effectiveness of our proposed method.
-This method can construct multi-turn dialogues based on medical, financial, and legal QAs, thereby alleviating the dialogue scarcity in other application domains.
-  • To better assess the quality of SmileChat, we collect a real-life counseling dataset with 82 counseling dialogues to build an authentic test set, PsyTest, which contains 4045 utterances spoken by supporters.
-Both automatic and human evaluations demonstrate that SmileChat improves the performance of the dialogue system in mental health.
-  • We release SmileChat, which comprises 55,165 Chinese multi-turn dialogues with an average of 10.4 turns.
-Additionally, we make our dialogue model, MeChat, and real-life test set, PsyTest, publicly available.
+  • To better assess the quality of SMILECHAT, we collect small-scale real-life counseling data with 50 anonymized counseling sessions to build a real-life test set, PsyTest.
+Automatic and human evaluations on the small-scale real-life test set confirm that our proposed dataset is high-quality.
+  • We release SMILECHAT, which consists of 55165 Chinese multi-turn dialogues. Our dialogue model, MeChat, and real-life test set, PsyTest, are publicly available.
+We highlight that this method can also construct multi-turn dialogues based on medical, financial, and legal QAs, thereby alleviating the dialogue scarcity in other application domains.
 2 Related Work
@@
 ChatGPT has proven to be a powerful AI tool for various NLP tasks since its release.
-Currently, it is being utilized in several domains, such as conversational AI (Alessa and Al-Khalifa, 2023; Köpf et al., 2023; Chen et al., 2023), education (Küchemann et al., 2023; Eshghie and Eshghie, 2023), code programming (Dong et al., 2023; Yetiştiren et al., 2023) and healthcare (Zhao et al., 2023; Yang et al., 2023).
+Currently, it is being utilized in several domains, such as conversational AI (Alessa and Al-Khalifa, 2023; Köpf et al., 2023; Chen et al., 2023), education (Küchemann et al., 2023; Eshghie and Eshghie, 2023), code programming (Dong et al., 2023; Yetiştiren et al., 2023), and healthcare (Zhao et al., 2023; Yang et al., 2023).
 Furthermore, ChatGPT’s efficiency and cost-effectiveness have been well-documented, making it competitive to human annotators (Gilardi et al., 2023; Zhu et al., 2023) even in zero-shot accuracy tasks.
@@
 (2023) have suggested using ChatGPT for task planning when receiving user inquiries, selecting appropriate models based on function descriptions from Hugging Face, executing each subtask using the chosen AI model, and summarizing the response based on the execution’s outcomes.
-In summary, ChatGPT has already demonstrated its enormous potential as an intelligent pipeline tool that can significantly advance NLP development, despite having only a restricted API available for researchers.
+In summary, ChatGPT has already demonstrated its enormous potential as an intelligent pipeline tool that can significantly advance NLP development despite having only a restricted API available for researchers.
 2.2 Datasets for Mental Health Support
@@
 Liu et al.
-(2021) first define the emotional support conversation task and then, via crowd-sourcing, construct ESConv, an emotional support conversation dataset containing 1053 dialogues with rich support strategies.
-However, the collection of ESConv requires high cost and time yet leads to a small-scale dialogue dataset.
+(2021) first defined the emotional support conversation task and then, via crowdsourcing, constructed ESConv, an emotional support conversation dataset containing 1053 dialogues with rich support strategies.
+However, the data collection of ESConv requires high cost and time yet leads to a small-scale dialogue dataset.
 To this end, Zheng et al.
-(2022) present an approach for augmenting data scale with informative dialogue posts and then constructing AugESC, a model-synthesized dataset with 102k dialogues.
+(2022) presented an approach for augmenting data scale with informative dialogue posts and then constructing AugESC, a model-synthesized dataset with 102k dialogues.
 The previous two datasets are limited to English.
 To facilitate the research in Chinese, hence Sun et al.
-(2021) crawl QA posts in a public mental health support platform compiling PsyQA.
+(2021) crawled QA posts on a public mental health support platform and made the PsyQA dataset publicly available.
 3 Method
-PsyQA [^2] https://www.xinli001.com/qa, a high-quality Chinese dialogue dataset related to mental health support in the form of one question mapping to multiple answers, has gone through a data anonymization process.
-Our dataset creation pipeline based on PsyQA comprises three main stages: (1) data preprocessing, (2) dialogue generation, and (3) dialogue retention.
-3.1 Data Preprocessing
-Considering the distinction between QAs in PsyQA and multi-turn dialogues and the context window limitation of 4096 tokens in ChatGPT, we recommend performing data preprocessing for PsyQA.
+PsyQA [^2] https://www.xinli001.com/qa, an open-source and high-quality Chinese single-turn dialogue dataset focused on mental health support, features one question mapped to multiple answers.
+Our dataset creation pipeline, based on PsyQA, includes three main stages: (1) data preparation, (2) task definition, and (3) prompt design.
+3.1 Data Preparation
+Considering the distinction between QA within PsyQA and multi-turn dialogues, along with the context window limitation of 4096 tokens in ChatGPT [^3] The model we use is gpt-3.5-turbo-0613., we propose to perform data preprocessing for PsyQA.
 This process involves wording cleaning and length truncation.
 Wording Cleaning
-This work aims to construct a large-scale, diverse, and high-quality multi-turn conversation corpus using the proposed smile method based on PsyQA.
-While QA can be considered a single-turn conversation between a real help-seeker and a supporter, there are differences in wording compared to actual multi-turn conversations.
-For instance, the term "楼主" (literally meaning "thread starter") frequently appears in QA but is rarely used in conversation.
+This work aims to construct a large-scale, lifelike, diverse, and high-quality multi-turn conversation corpus using the proposed SMILE method based on PsyQA.
+While QA can be considered a single-turn conversation between a real help-seeker [^4] The terms help-seeker and client are interchangeable.
+and a supporter [^5] The terms supporter and counselor are interchangeable., there are differences in wording compared to actual multi-turn conversations.
+{CJK*}UTF8gbsnFor instance, the term "楼主" (literally meaning "thread starter") frequently appears in QA but is rarely used in conversation.
 Therefore, we propose a two-stage process to clean the wording in PsyQA, mitigating linguistic discrepancies before rewriting QA into multi-turn conversations.
 This process includes both automatic and manual cleaning procedures.
-For detailed content, see Appendix A.
+Please refer to Appendix A for a detailed process.
 Length Truncation
-After conducting a statistical analysis of the PsyQA dataset released by Sun et al.
-(2021), we find that 757 QAs had a total length exceeding 1800 characters.
-Additionally, we identify 9 QAs in which the total discourse length exceeds 4000 characters.
-Furthermore, the model gpt-3.5-turbo [^3] The model we use is gpt-3.5-turbo-0613.
-has a maximum context length of 4096 tokens.
+After a statistical analysis of the PsyQA dataset, we find that 757 QAs have a total length exceeding 1800 characters.
+Also, we identify 9 QAs, the total discourse length exceeding 4000 characters.
+However, the model gpt-3.5-turbo has a maximum context length of 4096 tokens.
 To ensure reliable and smooth rewrites, we limit the length of the QA pairs, maximizing the number of rewritten dialogue turns.
-Specifically, we truncate the length of the QA pairs at 1800 characters and truncate any excess text.
-This control ensures that the generated text is limited to approximately 2,000 tokens.
+Specifically, we truncate the length of the QA pairs at 1800 Chinese characters and truncate any excess text.
+This control ensures that the generated text is limited to approximately 2000 tokens.
 It is worth noting that the PsyQA data used in this study undergo a data preprocessing process.
-3.2 Dialogue Generation
-3.2.1 Task Formulation
-Considering the remarkable capabilities of ChatGPT and its cost-effective, plug-and-play characteristics, we consider ChatGPT to be a suitable choice for this task compared to other alternative models.
-Our study is designed to prompt ChatGPT to rewrite a single-turn dialogue into a multi-turn one.
-Generally, text generation can be formulated as follows:
-$P_{\mathcal{M}_{\theta}}(Y|x)=\prod_{t=1}^{L}P_{\mathcal{M}_{\theta}}(y_{t}|Y_{<t},x)$ (1)
-where $\mathcal{M}$ denotes a large language model with parameters $\theta$.
-$Y$ represents the output of the model generation, while $L$ signifies the variable length of the output.
-The symbol $y_{t}$ represents the $t$-th token generated by the model, and $x$ denotes the conditional model input.
-3.2.2 Prompt Design
+3.2 Task Definition
+Let us first denote the input $x$ as a sequence $\{x_{1},x_{2},...,x_{n}\}$, and the output $y$ as a sequence $\{y_{1},y_{2},...,y_{m}\}$.
+The generation process of the language model can be expressed as the conditional probability distribution $p(y|x)$, which represents the probability of generating output $y$ given the input $x$.
+Therefore, text generation via the large language model can be formulated as follows:
+$p(y|x)=\prod_{t=1}^{m}p(y_{t}|y_{<t},x)$ (1)
+where $y_{t}$ represents the $t$-th token generated by the model. Because $x$ is our main focus in this paper, we will demonstrate the details of the prompt design next.
+3.3 Prompt Design
+In this section, we mainly focus on describing prompt design.
+In order to provide a clearer understanding of our method in a more controllable setting and elucidate the superiority of introducing single-turn dialogues as a reference, we first design two baseline prompts for comparison.
+3.3.1 standard Prompt
+As its name suggests, the standard prompt does not contain any single-turn dialogues or specific dialogue topics and instead uses only the most straightforward prompt to generate multi-turn dialogues.
+The standard prompt is illustrated in Figure 9 in Appendix C.
+The input in Equation 1 is $x=I$, where $I$ represents the standard prompt.
+We simplify the method name as standard and consider this method as our baseline.
+3.3.2 standardT Prompt
+Intuitively, feeding a single, fixed prompt into a large language model often generates text with low diversity.
+Therefore, we provide a specific dialogue topic for the standardT prompt.
+The input in Equation 1 is $x=(I,T)$, where $T$ represents the dialogue topic chosen in uniform sampling in the topic set.
+We simplify the method name to standardT and adopt this method as our baseline, as illustrated in Figure 10 in Appendix C.
+3.3.3 SMILE Prompt
+Our paper aims to highlight the superiority of introducing single-turn dialogues when generating dialogues.
+Our proposed method, the SMILE method, instructs the LLM to rewrite single-turn dialogues into multi-turn ones.
+Figure 2 depicts the concrete prompt details.
+The input in Equation 1 is $x=(I,T,D)$, where $T$ and $D$ represent the implicit dialogue topic hidden in the QA and single-turn dialogue, respectively.
 Figure 2: The SMILE method used to generate dialogues for mental health support.
-To establish some conventions regarding specific terms for prompt design, we define the "help-seeker" and "supporter" as "求助者" and "支持者" in Chinese, respectively.
-A single-turn dialogue is defined as "求助者:$u^{H}$持者:$u^{S}$", where $u^{H}$ and $u^{S}$ represent the utterances of the help-seeker and supporter, respectively, and $H$ and $S$ refer to the help-seeker and supporter.
-In this section, we mainly focus on detailing our methodology. Furthermore, we are building two baseline methods for comparison.
-Standard Prompt
-As its name suggests, the standard prompt does not contain any single-turn dialogues and instead uses only the simplest prompt to generate multi-turn dialogues.
-The specifics of the standard prompt are illustrated in Figure 8 in Appendix C.
-The conditional model input in Equation 1 is $x:=(\mathcal{I})$, where $\mathcal{I}$ represents the standard prompt.
-We simplify the method name as standard and consider this method as our baseline.
-Standard Prompt with a Dialogue Topic
-Intuitively, feeding a single, fixed prompt into a large language model often results in the generation of low diversity.
-To address this issue, we collaborate with three professional counselors, refer to existing literature (Rickwood et al., 2007; Pedrelli et al., 2015), and ultimately compile a comprehensive set of dialogue topics.
-This set comprises 56 distinct types, each accompanied by its corresponding explanation.
-For more details, please refer to Appendix K.
-The conditional model input in Equation 1 is $x:=(\mathcal{I,T})$, where $\mathcal{I}$ represents the dialogue topic chosen in uniform sampling.
-We simplify the method name to stand.${}_{\mathrm{DT}}$ and adopt this method as our baseline, as illustrated in Figure 8 in Appendix C.
-smile Method
-Our proposed method, referred to as the smile method, instructs the ChatGPT to rewrite single-turn dialogues into multi-turn ones.
-Figure 2 depicts the specific prompt template.
-The conditional model input in Equation 1 is $x:=(\mathcal{I,T,D})$, where $\mathcal{T}$ and $\mathcal{D}$ represent the dialogue topics hidden in the QA and single-turn dialogue, respectively.
-3.3 Dialogue Retention
-ChatGPT may exhibit potential instability during the dialogue generation process, but this occurs very infrequently.
-Therefore, we employ an automatic filtering mechanism to exclude dialogues that fail to meet our specified requirements, which encompass dialogue format and dialogue turns as detailed in Appendix B.
 4 Language Transformation
 4.1 Experimental Setup
+Dialogue Topic Collection
+To address the issue of monotonous generation in the standard prompt, we propose to enhance diversity by injecting a dialogue topic into the standardT prompt.
+Consequently, we have collaborated with three professional counselors, reviewed existing literature (Rickwood et al., 2007; Pedrelli et al., 2015), and ultimately compiled a comprehensive set of dialogue topics.
+This set includes 60 distinct types of dialogue topics, each with a detailed explanation.
+For more details, please refer to Appendix G.
 QA Sampling
-To ensure a fair comparison and prevent duplicate instances of the same question with different answers, we first randomly select 500 non-duplicate questions from the first 5,000 QAs in PsyQA.
-We then randomly choose one answer to serve as the corresponding question’s response.
+To ensure a fair comparison among the three methods we propose and prevent duplicate instances of the same question with different answers, we first randomly select 500 non-repetitive questions from the first 5000 QAs in PsyQA.
+We then randomly choose one answer to serve as the corresponding response.
 The data samples obtained are employed as seed dialogues, which are subsequently restructured into multi-turn conversations via ChatGPT.
-We name this 500 sampled data as PsyQA* to distinguish it from the whole dataset.
+We name this 500 sampled data as PsyQA* to distinguish it from the original dataset.
 Hyperparameters
-On the whole, we present three prompt methods in this paper.
+Overall, we present three prompt methods in this paper.
 For each prompt method, we instruct ChatGPT to generate 500 dialogues to study language transformation and dialogue diversity, thereby validating the feasibility and effectiveness of our proposed SMILE method, respectively.
-To enhance the diversity of the generated dialogues, we set ChatGPT’s hyperparameters during text generation to the officially recommended default values, with $\tau$ (temperature) = 1.0 and $p$ (top_p) = 1.0.
+To enhance the diversity of the generated dialogues, we set hyperparameters during text generation to the officially recommended default values, with temperature = 1.0 and top_p = 1.0.
 Dialogue Filtering
-[…]
-Text Representation
-Text representation is used for analyzing language transformation and semantic diversity.
-A dialogue between a help-seeker and a supporter is represented as $d=\{u_{0}^{H},u_{1}^{S},u_{2}^{H},...,u_{t}^{X},...,u_{n-1}^{H},u_{n}^{S}\}$, where $u_{t}^{X}$ represents the utterance of the $t$-th turn spoken by either the help-seeker or supporter.
-A string of a dialogue can be denoted as $d_{s}=[u_{0}^{H};u_{1}^{S};u_{2}^{H};...;u_{n-1}^{H};u_{n}^{S}]$, where [;][;][ ; ] denotes the operation of textual concatenation.
-To obtain the text embedding of a dialogue, we employ OpenAI’s model text-embedding-ada-002444https://platform.openai.com/docs/guides/embeddings, which accepts a maximum context length of 8191.
+ChatGPT may exhibit potential instability during the text generation process.
+Therefore, as detailed in Appendix B, we employ an automatic filtering mechanism to exclude dialogues that fail to meet our specified requirements, which encompass dialogue format and dialogue turns.
+When a dialogue falls short of these criteria, we prompt ChatGPT to generate the dialogue anew until it aligns with our specified requirements only in the initial 500 dialogues.
+This measure, therefore, ensures that each method produces an equal number of 500 dialogues for preliminary analysis.
+Text Embedding
+A multi-turn dialogue between a help-seeker and a supporter is represented as
+$d=\{u_{1},r_{1},...,u_{i},r_{i},...,u_{n},r_{n}\}$ (2)
+where $u_{i}$ and $r_{i}$ represents the utterances of the $i$-th turn spoken by the help-seeker and supporter, respectively.
+A string of dialogue without any speaker role tokens can be denoted as $d_{s}=[u_{1};r_{1};u_{2};r_{2};...;u_{n};r_{n}]$, where [;][;][ ; ] denotes the operation of textual concatenation.
+To obtain the text embedding of a dialogue, we employ OpenAI’s model text-embedding-ada-002666https://platform.openai.com/docs/guides/embeddings, which accepts a maximum context length of 8191.
 Each dialogue is first preprocessed into a single string without any speaker tokens and is then mapped to a 1536-dimensional vector.
@@
 4.2 Analysis
-[…]
-Figure 3 presents the distribution of dialogue transformation among three methods.
-According to our analysis, we conclude that single-turn dialogues can be successfully rewritten into multi-turn dialogues.
-Figure 3: Distribution of dialogue transformation among three methods. The line $x=0.9312$ represents the boundary of $\mu-3\sigma$ in the smile method.
-5 Dialogue Diversity
-[…]
-Figure 4: Pairwise dialogue cosine similarity among four settings: our proposed three methods and a reference point using sampled data from PsyQA.
-5.1 Semantic Features
-[…]
-To calculate the pairwise cosine similarity between every two distinct dialogues, this produces $\binom{500}{2}$ pairs of dialogues and their corresponding cosine values in each prompt method, including the reference point of sampled data from PsyQA.
-We present the results in Figure 4, which demonstrates that the median of the smile method is significantly lower than those of the baseline methods.
-The smile method exhibits the most extensive semantic diversity, aligning closely with the sampled dialogue of PsyQA.
-However, it’s worth noting that the median of the smile method is larger than that of PsyQA*.
-The reason behind this may be the introduction of token distribution from ChatGPT.
-Category Total Help-seeker Supporter
-# Dialogues 55165 - -
-# Utterances 1833856 693756 1140100
-# Avg. turns per dialogue 10.4 - -
-Avg. utterances per dialogue 33.2 12.6 20.7
-Avg. length per utterance 27.9 26.1 28.9
-Table 1: Statistics of the dialogue dataset, SmileChat.
-6 Expert Evaluation
-We elaborate on data quality analysis from two main professional perspectives: (1) the evaluation of dialogue quality by three professional counselors, and (2) alignment with the supporters’ strategy flow patterns in real-life counseling conversations (For brevity, we include this section in Appendix G.).
-To further validate that the dialogue quality generated using the SMILE method is superior to that produced by the two baseline methods, we propose to recruit three professional psychological counselors and randomly sample 20 dialogues for each method.
-Each set comprises three dialogues generated from different methods.
-Three counselors assign scores (3, 2, and 1, where 3 is the best, and 1 is the worst) based on three metrics: assessing (Q1) the overall quality of the conversation, (Q2) the authenticity of the user’s stated issues, and (Q3) the professionalism of the counselor’s responses.
-We present the results of an expert evaluation of dialogue quality in 20 sets of dialogues in Table 2.
-Each value in the table represents the average score from three professional counselors, and results show that the dialogue quality generated with the SMILE method is of higher quality than that generated by the two baseline methods.
-The intraclass correlation coefficients (James et al., 1984; Koo and Li, 2016), including nine coefficients for the three metrics, are all greater than 0.8, indicating almost perfect agreement.
-Method Q1 Q2 Q3
-standard 1.95 2.12 1.21
-stand.${}_{\mathrm{DT}}$ 1.25 1.2 1.82
-smile 2.8 2.68 2.97
-Table 2: Evaluation of dialogue quality by three professional counselors.
-7 SmileChat Dataset
-Through the analysis of language transformation, dialogue diversity and expert evaluation, we conclude that the proposed method can generate a diverse and high-quality chat dataset.
-Therefore, we utilize the smile method to guide ChatGPT in generating all multi-turn conversations based on PsyQA one round, leading to a large-scale dialogue dataset.
-7.1 Data Statistics
-To ensure data quality, we impose stricter requirements on dialogue turns, retaining only dialogues with at least 5 turns.
-Thus, we compile a collection of 55,165 conversations, SmileChat.
-Table 1 presents the statistics of the collected corpus.
-7.2 Dialogue Exemplars
-Multi-turn dialogue examples generated with the standard and stand.${}_{\mathrm{DT}}$ methods are illustrated in Figure 13 and 15.
-Further, a dialogue generated by the smile method is shown in Figure 17.
-8 Dialogue System
-We aim to build a high-quality multi-turn chat dataset for mental health. Therefore, we also analyze the dialogue quality based on the performance of the dialogue system trained with SmileChat.
-8.1 Task Formulation
-Our collected dataset can be represented as $D=\{d_{1},d_{2},...,d_{n}\}$, where each $d_{j}$ represents a single multi-turn dialogue.
-To train a dialogue system for mental health, we need to split each dialogue into several training sessions. Specifically, a sampled $t$-turn dialogue session can be represented as follows:
-$d_{t}=\left\{u^{H}_{0},u^{S}_{1},u^{H}_{2},...,u^{H}_{t-1},u^{S}_{t}\right\}\sim D$ (3)
-We build a dialogue model that can predict the supporter’s utterance $u^{S}_{t}$ based on the dialogue history $h_{t}=\left\{u^{H}_{0},u^{S}_{1},u^{H}_{2},...,u^{H}_{t-1}\right\}$.
-Our objective is to maximize the likelihood probability as follows:
-[…]
-where $L$ is the sequence length of $u^{S}_{t}$.
-8.2 Experimental Setup
-Baseline Model
-To validate the dialogue quality of our collected dataset, we conduct a fine-tuning experiment on ChatGLM2-6B (Zeng et al., 2023).
-Moreover, we also add two prompt engineering baselines (Prompt A and B, for more details, see Appendix I).
-Parameter-efficient Fine-tuning
-To preserve the original capabilities of the model while adapting to downstream dialogue tasks and reducing computational costs, we employ Low-Rank Adaptation (LoRA, (Hu et al., 2021)) on all linear layers in the ChatGLM2-6B model for efficient fine-tuning.
-Hyperparameters
-We present the hyperparameters for constructing a dialogue model for mental health in Table 10. For additional details on the training and generation processes, please refer to Appendix H.
-Category Total Help-seeker Supporter
-# Dialogues 82 - -
-# Utterances 7672 3627 4045
-# Avg. turns per dialogue 74.8 - -
-Avg. utterances per dialogue 93.6 44.2 49.3
-Avg. length per utterance 24.6 29.7 20.1
-Table 3: Statistics of test set, PsyTest.
-Methods PPL ($\Downarrow$) METEOR ($\Uparrow$) BLEU-1 ($\Uparrow$) BLEU-2 ($\Uparrow$) BLEU-3 ($\Uparrow$) Rouge-L ($\Uparrow$) D-1 ($\Uparrow$) D-2 ($\Uparrow$) D-3 ($\Uparrow$)
-Baseline 3.38 ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$
-Baseline + Prompt A 3.43 ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$
-Baseline + Prompt B 3.99 ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$
-Fine-tuned 1.50 ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$ ${}^{\dagger}$${}^{\ddagger}$
-Table 4: Results of automatic evaluation in our compiled real-life test set.
-$\dagger$ denotes the tokenizer we use is THUDM/chatglm2-6b, while $\ddagger$ denotes the tokenizer we use is hfl/chinese-roberta-wwm-ext-large.
-8.3 Evaluation
-To better understand and assess the dialogue quality of SmileChat dataset, we propose to utilize real-life multi-turn counseling conversations.
-We develop an online mental health support platform that enables professional counselors to offer each client a free text-based counseling service, lasting approximately 50 minutes each time.
-This duration is widely recognized as a standard time setting in psychological counseling.
-Therefore, we compile a collection of 82 high-quality real-life counseling dialogues with data desensitization for model evaluation.
-We name this test set PsyTest, which contains 4045 utterances spoken by supporters, as shown in Table 3.
-8.3.1 Automatic Evaluation
-Metrics
-To conduct automatic evaluation, the evaluation metrics we use consist of Perplexity (PPL) (Jelinek et al., 1977), METEOR (Banerjee and Lavie, 2005), BLEU-1/2/3 (Papineni et al., 2002), Rouge-L (Lin, 2004), and Distinct-1/2/3 (D-1/2/3) (Li et al., 2016).
-Results
-The results of the automatic evaluation, including 9 metrics, are presented in Table 4.
-Notably, the evaluated dialogues are based on real-world counseling data rather than generated dialogues, which excludes the influence stemming from ChatGPT.
-All automatic evaluation metrics we use indicate improved performance.
-Our results show that the model trained with SmileChat is effective and practical.
-Consequently, the automatic evaluation demonstrates that our collected dataset is of high quality.
-8.3.2 Human Evaluation
-Metrics
-We conduct a pairwise human evaluation to study the model performance trained with our proposed dialogue corpus.
-Initially, we randomly sample 100 multi-turn dialogue sessions, each comprising a minimum of 9 turns selected from PsyTest.
-Each dialogue session concludes with the utterance spoken by the counselor (also referred to as the supporter).
-Subsequently, we obtain 100 generated responses from the MeChat model.
-Three professional counselors are then presented with a dialogue history and three randomly shuffled responses (baseline, fine-tuned, ground truth).
-They are tasked with selecting the optimal response for the dialogue history, considering aspects such as naturalness, human-like qualities, correctness, and appropriateness.
-The evaluation is conducted based on the ethical principles of psychologists and the code of conduct (Association et al., 2016).
-Figure 5: Pairwise human evaluation results, comparing the fine-tuned model to the baseline and ground truth, are reported.
-We present the win and lose rates of each compared pair in 100 randomly sampled real-life dialogue sessions.
-Fleiss’ kappa (Fleiss et al., 1981) is used to measure the inter-rater agreement, and all values fall within moderate agreement with $0.5\leq\kappa\leq 0.6$.
-Results
-We employ majority voting to reach final decisions among three professional counselors.
-As depicted in Figure 5, the model trained with SmileChat demonstrates a significant performance improvement compared to the baseline model.
-Moreover, the responses generated by MeChat surpass the golden response to some extent in 36% of all dialogue sessions.
-Case Study
-We present a case study, as shown in Figure 6.
-First, during counseling conversations, the text should not be too long for the help-seeker to understand.
-Second, at the beginning of a conversation, a supporter generally will not challenge the help-seeker; instead, they will provide support.
-Third, response 1 is more human-like and natural.
-Therefore, response 1 is better than response 2 overall.
-Figure 6: Case study. Counseling strategies used in the two responses are presented in parentheses. Strategies in green are supporting, while those in red are challenging.
-9 Conclusion
-This paper introduces smile, a simple yet effective solution for addressing the scarcity of multi-turn conversations in mental health.
-Through language transformation and diversity analysis, we confirm the feasibility and effectiveness of our approach, demonstrating that our method is transferable.
-We also conduct expert evaluations, demonstrating that the smile method exhibits high superiority.
-Thus, the proposed method enables the automatic creation of a large-scale, diverse, and high-quality dialogue corpus, SmileChat, consisting of 55,165 dialogues with an average of 10.4 turns.
-Both automatic and human evaluations using the PsyTest dataset, consisting of 82 real-life counseling dialogues, demonstrate that SmileChat significantly improves dialogue system performance in mental health.
-With the release of multi-turn dialogues (SmileChat), our dialogue model (MeChat), and an authentic test set (PsyTest), we contribute valuable resources to the research community.
-Limitations
-We release a large-scale, diverse, and high-quality multi-turn conversational dataset for mental health support, generated by rewriting single-turn conversations into multi-turn conversations using ChatGPT.
-Consequently, the dataset unavoidably incorporates ChatGPT’s model knowledge.
-Furthermore, we discuss how ChatGPT does not fully utilize the rich vocabulary and content of single-turn conversations, as reflected in the distinct-$n$ metric.
-Additionally, there is inevitably a gap in counseling practicality between models trained on this synthetic dataset and those trained on 50-minute counseling conversations, which will be our future work.
-While dialogues generated by the three methods still exhibit significant gaps when compared to the Xingling (Oracle) dataset with respect to the flow patterns of supporters’ strategies, there exists a large space for us to improve in our future work.
-Furthermore, there is currently no comprehensive dataset available for evaluating the conversational safety of mental health support models.
-Therefore, we have identified this as a limitation of our paper and intend to address it in future work.
-Ethical Considerations
-We present a detailed illustration for the section on ethical considerations, which includes Data Usage Agreement, Usage Policy, and Suggestions for Real-life Deployment.
-Furthermore, the section on suggestions for real-life deployment highlights four main aspects: Anonymity and Consent, Handling Sensitive Topics, Ensuring Client Confidentiality, and Preventing Misinformation.
-Data Usage Agreement
-PsyQA is a single-turn dialogue dataset collected from an online mental health support platform.
-Specifically, help-seekers submit a post containing their mental health states and issues, and many professional counselors write down their responses to user questions to assist help-seekers.
-Therefore, PsyQA is a high-quality Chinese dialogue related to mental health support in the form of one question mapping to multiple answers.
-Following the data copyright guidelines formulated by PsyQA (Sun et al., 2021), we release the multi-turn dialogue corpus publicly available for research purposes only.
-If researchers wish to reproduce the multi-turn dialogues using PsyQA, they must sign an agreement with the original data owner.
-Accordingly, we release our datasets and models for research purposes, thus facilitating further advancement in the academic community.
-Usage Policy
-Our chatbot is trained with the SmileChat dataset end-to-end.
-As illustrated in our paper, SmileChat still exhibits a significant gap in the flow patterns of counseling strategies compared to real-life counseling dialogues.
-Access to SmileChat and Mechat should be strictly limited to research purposes and should NOT be utilized for real-world deployment, commercial purposes, or any other use beyond academic research.
-Researchers using SmileChat and MeChat should be aware of the dataset and model limitations and make efforts to acknowledge and/or mitigate them to the best of their ability.
-Suggestions for Real-life Deployment
-When deploying our chatbot on the server for research purposes, we recommend implementing safeguards to protect help-seekers and, to a large extent, mitigate potential physical risks.
-Anonymity and Consent
-We strongly recommend that researchers allow users to interact with the chatbot anonymously, especially when discussing sensitive issues.
-Additionally, researchers should clearly communicate the level of anonymity and obtain informed consent from users before collecting any personal information.
-Handling Sensitive Topics
-In real-life human-machine interactions, we strongly recommend constructing a classifier to identify user intent.
-When users discuss sensitive topics, if detected by the intent classifier, a "human-in-the-loop" system allows a human moderator or healthcare professional to intervene in complex or sensitive situations.
-Furthermore, we recommend ensuring clear escalation paths for users who require immediate human assistance, such as in cases of suicide and self-injury.
-Ensuring Client Confidentiality
-Researchers should clearly disclose to users that they are interacting with a chatbot, not a human, setting appropriate expectations for the system’s capabilities and limitations.
-Researchers should provide clear information on how data is handled, stored, and for what purposes.
-Misinformation Prevention
-Researchers should implement fact-checking mechanisms to verify the accuracy of information provided by the chatbot.
-Researchers should clearly communicate the limitations of the chatbot’s knowledge and advise users to consult a human professional for critical decisions.
-Appendix A Details of Data Cleaning
-A.1 Automatic Cleaning
-We employ a sequential data cleaning pipeline to swiftly replace words that are unsuitable to the conversation scenario.
-For example, both "楼主你" (literally thread starter you) and "楼主" (literally thread starter) should be replaced with "你" (you).
-However, it is necessary to perform the former replacement to avoid the repetition of "你" and the resulting "你你" (you-you).
-A.1.1 Word List for Data Cleaning
-To avoid the repetition of "你" (you) and the resulting "你你" (you-you) , we suggest to conduct a sequential word replacing pipeline.
-Figure 7 shows the word list for data cleaning and corresponding order for automatic cleaning.
-Figure 7: Word list for automatic cleaning.
-A.2 Manual Cleaning
-Due to the specificity and complexity of language, manual cleaning remains an essential part of the process.
-To prevent virtual dialogue systems from exhibiting overly frequent anthropomorphic behavior, we identify instances of the Chinese word for "hug" (抱抱) and manually delete sentence snippets containing this term.
-Appendix B Requirements for Dialogue Retention
-Here are two main requirements for dialogue filtering: data format and dialogue turns.
-B.1 Data Format
-We provide the requirements for data format as follows:
-  • 1.
-    The generated conversations do not start with "求助者:" or "支持者:".
-  • 2.
-[…]
-  • 3.
-    Each utterance in generated conversations does not start with "求助者:", "求助者:", "支持者:" or "支持者:".
-  • 4.
-    The last utterance in generated conversations contains an English sentence.
-B.2 Dialogue Turns
-Conversations comprising fewer than 5 turns will be discarded.
-Appendix C Method
-For brevity, we present the standard prompt and the standard prompt with a dialogue topic in Figure 8.
-Figure 8: The standard prompt and the standard prompt with a dialogue topic.
-The two are separated by using the mark of "/" to differentiate between them.
-The key difference lies in whether a dialogue topic is introduced.
-Appendix D Mechanism of Language Transformation
 We propose two hypotheses: (1) When a single-turn dialogue is rewritten into a multi-turn dialogue, the similarity between the two is high (Attract).
 (2) When ChatGPT generates a multi-turn dialogue without introducing a single-turn dialogue, the similarity between the generated dialogue and a randomly selected single-turn dialogue is low (Repel).
-For better understanding, we present the mechanism of language transformation in Figure 9.
+The mechanism of our proposed hypotheses is presented in Figure 3.
+To assess the transformation feasibility of our SMILE method, we employ cosine similarity to calculate the transformation ratio.
+Specifically, a single-turn dialogue d is transformed into a multi-turn dialogue $\hat{d}$ using the SMILE method.
+A multi-turn dialogue generated by standard or standardT is denoted as $\bar{d}$.
+We can calculate cos($d$, $\hat{d}$) for the first hypothesis and cos($d$, $\bar{d}$) for the second hypothesis.
+We utilize the text embeddings obtained in Equation 3.
 Figure 3: Mechanism for language transformation.
-Appendix E Dialogue Topics Annotation
-In this paper, to label the dialogue topics of generated dialogues, the hyperparameters of ChatGPT API during generation we used are set to the officially recommended default values, where temperature $\tau=1.0$ and $p=1.0$.
-Figure 10 shows the pipeline of dialogue topics annotation.
-Figure 10: Pipeline of Dialogue Topics Annotation, where the content in bold is a placeholder.
-Appendix F Dialogue Diversity
-Metrics
+Figure 4: Distribution of dialogue transformation among three methods. The line $x=0.9312$ represents the boundary of $\mu-3\sigma$ in the SMILE method.
+Figure 4 presents the distribution of dialogue transformation among three methods.
+Our analysis concludes that single-turn dialogues can be successfully rewritten into multi-turn dialogues, ensuring that the dialogue generated by the proposed method is lifelike, rather than purely in the style of machine-generated text from ChatGPT.
+Method
 # Unique
@@
 Unigrams
-Distinct-1 ($\Uparrow$)
+Distinct-1
 # Unique
@@
 Bigrams
-Distinct-2 ()($\Uparrow$)( ⇑ )
+Distinct-2
 # Unique
@@
 Trigrams
-Distinct-3 ()($\Uparrow$)( ⇑ )
+Distinct-3
 PsyQA* 11785 203306 0.058 120049 202806 0.592 182942 202306 0.904
 standard 4174 153536 0.027 32281 153036 0.211 72340 152536 0.474
-stand.${}_{\mathrm{DT}}$ 6032 175319 0.034 52141 174819 0.298 105921 174319 0.608
+standardT 6032 175319 0.034 52141 174819 0.298 105921 174319 0.608
 SMILE 10447 254585 0.041 111662 254085 0.439 196367 253585 0.774
 Table 1: Statistics of 500 conversations in each prompt method, including PsyQA*.
-F.1 Lexical Features
-For lexical analysis, we utilize the ChatGLM2-6B tokenizer [^5] https://huggingface.co/THUDM/chatglm2-6b to tokenize the dialogue.
+5 Dialogue Diversity
+To demonstrate the effectiveness of the SMILE method, we mainly focus on three aspects of diversity: lexical features, semantic features, and dialogue topics.
+5.1 Lexical Features
+For lexical analysis, we utilize the popular Chinese ChatGLM2-6B [^7] https://huggingface.co/THUDM/chatglm2-6b tokenizer that is widely used in NLP to tokenize the dialogue.
 To measure the lexical features, we adopt distinct-$n$ ($n=1,2,3$) metrics (Li et al., 2016), which are widely used for measuring the diversity of dialogue datasets.
 Each dialogue is first preprocessed into a single string without any speaker tokens.
-We provide statistics for 500 dialogues per prompt method, with PsyQA* serving as the reference point, detailed in Table 5.
-Our proposed smile method results in rich vocabularies, with a significantly higher number of unique unigrams, bigrams and trigrams compared to the baseline methods.
+We provide statistics for 500 dialogues per prompt method, as presented in Table 1.
+Our proposed SMILE method results in rich vocabularies, with a significantly higher number of unique unigrams, bigrams, and trigrams than the two baseline methods.
 Specifically, a simple and fixed prompt tends to produce monotonous output.
-When dialogue topics are incorporated into a single, fixed prompt, the model’s output demonstrates substantial diversification.
-Furthermore, the smile method outperforms the baseline methods in terms of Distinct-1, Distinct-2 and Distinct-3.
-F.2 Dialogue Topics
+The model output demonstrates substantial diversification when dialogue topics are incorporated into a single, fixed prompt.
+Furthermore, the SMILE method outperforms the baseline methods in terms of Distinct-1, Distinct-2, and Distinct-3.
+Figure 5: Pairwise dialogue cosine similarity among four settings: our proposed three methods and a reference point using sampled data from PsyQA.
+5.2 Semantic Features
+To measure the semantic diversity in a dialogue dataset, we propose to calculate the cosine similarity between every pair of different dialogues.
+This involves computing the pairwise cosine similarity for each pair of distinct dialogues, resulting in $\binom{500}{2}$ pairs and their corresponding cosine values, as described in Equation 3.
+We present the results in Figure 5, which demonstrates that the median of the SMILE method is significantly lower than those of the baseline methods.
+The SMILE method exhibits the most extensive semantic diversity, aligning closely with the sampled dialogue of PsyQA.
+However, it is worth noting that the median of the SMILE method is more significant than that of PsyQA*.
+The reason behind this could be the introduction of token distribution from ChatGPT.
+5.3 Dialogue Topics
 To measure the diversity of dialogue topics in a dialogue dataset, we utilize information entropy to measure the diversity of topic distribution.
@@
 where $H(X)$ is the information entropy. $p(x_{i})$ is the probability of the occurrence of topic $x_{i}$.
-To obtain dialogue topics for each dialogue in each prompt method, we design a prompt provided with 56 distinct dialogue topics, as illustrated in Appendix K and Figure 10, to automatically label dialogue topics for each dialogue with GPT-4.
-The model we use is gpt-4-0613.
-We present the information entropy for each prompt method in Table 6, demonstrating that the dialogues generated using the smile method are substantially more diverse than those generated using the standard method and are compatible with the stand.${}_{\mathrm{DT}}$ method, which uniformly samples dialogue topics.
-Setting standard stand.${}_{\mathrm{DT}}$ smile
+To obtain dialogue topics for each dialogue in each prompt method, we design a prompt provided with 60 distinct dialogue topics, as illustrated in Appendix G and Figure 11, to automatically label dialogue topics for each dialogue with Qwen1.5-110B-Chat (qwen).
+We present the information entropy for each prompt method in Table 2, demonstrating that the dialogues generated using the SMILE method are substantially more diverse than those generated using the standard method and are compatible with the standardT method, which uniformly samples dialogue topics.
+Setting standard standardT SMILE
 Limited Topics 8.11 14.28 14.07
@@
 Table 2: Information entropy of dialogue topics.
-The setting of limited topics indicates that the dialogue topics in the statistics are confined to the 56 dialogue topics we provide.
-The setting of unlimited topics denotes that the dialogue topics not only belong to our provided 56 distinct topics but also include topics generated spontaneously by GPT-4.
-Appendix G Expert Evaluation
-To illustrate the patterns of supporters’ strategies, we compile statistics on the flow patterns of these strategies, including "Supporting", "Challenging" and "Other".
-G.1 Experimental Setup
-PsyQA (Sun et al., 2021) contains 4012 QAs with supporters’ strategies in the original dataset.
-Xinling (Li et al., 2023) is a real-life multi-turn counseling dataset containing 550 dialogues with supporters’ strategies in the original dataset.
-Three methods (standard, stand.${}_{\mathrm{DT}}$, and smile) each contain 500 dialogues, respectively.
-We use Xinling to train a classifier to predict the supporters’ strategies in dialogues generated with these three methods.
-The performance of the trained classifier is comparable to the results in the original paper.
-G.2 Results
-Supporters have three main strategies: "Supporting", "Challenging", and "Other", and we present the results in Table 7, 8 and 9.
-We highlight the gap greater the 0.2 in bold and the corresponding gap is in parentheses.
-Strategy Supporting
-Methods standard stand.DT SMILE PsyQA Xinling (Oracle)
-Stage 1 0.854(+0.101) 0.847(+0.094) 0.615(-0.138) 0.474(-0.279) 0.753
-Stage 2 0.547(-0.278) 0.520(-0.305) 0.432(-0.393) 0.312(-0.513) 0.825
-Stage 3 0.419(-0.367) 0.365(-0.421) 0.360(-0.426) 0.179(-0.607) 0.786
-Stage 4 0.468(-0.258) 0.387(-0.339) 0.390(-0.336) 0.107(-0.619) 0.726
-Stage 5 0.808(+0.382) 0.774(+0.348) 0.662(+0.236) 0.154(-0.272) 0.426
-Table 7: Pattern flow of "Supporting" strategy.
-Strategy Challenging
-Methods standard stand.DT SMILE PsyQA Xinling(Oracle)
-Stage 1 0.140(+0.084) 0.150(+0.094) 0.372(+0.316) 0.348(+0.292) 0.056
-Stage 2 0.452(+0.305) 0.478(+0.331) 0.565(+0.418) 0.598(+0.451) 0.147
-Stage 3 0.576(+0.384) 0.632(+0.440) 0.636(+0.444) 0.752(+0.560) 0.192
-Stage 4 0.516(+0.272) 0.603(+0.359) 0.586(+0.342) 0.847(+0.603) 0.244
-Stage 5 0.112(-0.127) 0.157(-0.082) 0.231(-0.008) 0.794(+0.555) 0.239
-Table 8: Pattern flow of "Challenging" strategy.
-Strategy Other
-Methods standard stand.DT SMILE PsyQA Xinling(Oracle)
-Stage 1 0.006(-0.185) 0.003(-0.188) 0.013(-0.178) 0.178(-0.013) 0.191
-Stage 2 0.001(-0.027) 0.002(-0.026) 0.003(-0.025) 0.090(+0.062) 0.028
-Stage 3 0.005(-0.017) 0.003(-0.019) 0.004(-0.018) 0.069(+0.047) 0.022
-Stage 4 0.016(-0.014) 0.010(-0.020) 0.024(-0.006) 0.046(+0.016) 0.03
-Stage 5 0.080(-0.255) 0.069(-0.266) 0.107(-0.228) 0.052(-0.283) 0.335
-Table 9: Pattern flow of "Other" strategy.
-G.3 Conclusion
-Dialogues generated by the three methods exhibit similar flow patterns of supporters’ strategies.
-Instructing ChatGPT to rewrite QAs into multi-turn dialogues will inevitably impact the strategy pattern in the generated dialogues.
-Dialogues generated by the three methods still exhibit significant gaps when compared to the Xingling (Oracle) dataset with respect to the flow patterns of supporters’ strategies.
-Appendix H Implementation Details
-Model Training.
-We implement ChatGLM2-6B with the Transformers library. Table 10 shows the parameters of parameter-efficient fine-tuning.
+5.4 SMILECHAT Dataset
+Through the comprehensive analysis of language transformation and dialogue diversity, we conclude that the proposed method can generate a lifelike and diverse chat dataset.
+Therefore, we utilize the SMILE method to guide ChatGPT in generating all multi-turn conversations based on PsyQA one round only, leading to a large-scale dialogue dataset.
+Category Total Help-seeker Supporter
+# Dialogues 55165 - -
+# Utterances 1833856 693756 1140100
+Turns per dialogue 5.7 - -
+Utterances per dialogue 33.2 12.6 20.7
+Avg. length per utterance 27.9 26.1 28.9
+Table 3: Data statistics of the dialogue dataset, SMILECHAT.
+Data Statistics
+To ensure data quality, we impose stricter requirements on dialogue turns, retaining only dialogues with at least five turns.
+Thus, we compile a collection of 55165 conversations, SMILECHAT.
+Table 3 presents the statistics of the collected corpus.
+Dialogue Exemplars
+Multi-turn dialogue examples generated with the standard and standardT methods are illustrated in Figures 18 and 20. Further, a dialogue generated by the SMILE method is shown in Figure 22.
+6 Dialogue System
+We aim to build a high-quality multi-turn chat dataset for mental health support.
+It is non-trivial to evaluate the quality of a dialogue dataset, which is often assessed indirectly by the dialogue system.
+Therefore, we need to train a dialogue system and analyze its performance.
+6.1 Mathematical Formulation
+In order to train a dialogue system for mental health support, the first step is to split each full dialogue $d\sim\mathcal{D}$ into several training sessions.
+Specifically, a sampled $t$-turn dialogue session can be represented as $d_{t}=\left\{u_{1},r_{1},u_{2},r_{2},\dots,u_{t},r_{t}\right\}\sim\mathcal{D}$.
+Thus, we build a dialogue model that can predict the supporter’s utterance $r_{t}$ based on the dialogue history $h_{t}=\left\{u_{1},r_{1},u_{2},r_{2},\dots,u_{t}\right\}$.
+Our objective is to use our synthetic conversation dataset $\mathcal{D}$ to fine-tune a large language model $\pi_{0}$ using supervised learning, i.e., maximum likelihood estimates:
+$J_{\mathrm{SFT}}(\theta)=\mathbb{E}_{(h_{t},r_{t})\sim\mathcal{D}}\left[\log\pi_{\theta}(r_{t}\mid h_{t})\right]$ (5)
+where $\pi_{\theta}$ is initialized from $\pi_{0}$.
+6.2 Experimental Setup
+Backbone Model
+To validate the dialogue quality of our collected dataset, we conduct a fine-tuning experiment on ChatGLM2-6B (Zeng et al., 2023).
+Training Data
+To meet the data format requirements for instruction-based fine-tuning, we split the dialogue into multiple sessions, with the supporter’s last utterance concluding each session.
+Additionally, we incorporate the system prompt (detailed in Appendix E) as a prefix to dialogue messages, following OpenAI’s data format.
+Parameter-efficient Fine-tuning
+To preserve the original capabilities of the model while adapting to downstream dialogue tasks and reducing computational costs, we employ Low-Rank Adaptation (LoRA, (Hu et al., 2021)) on all linear layers in the model for efficient fine-tuning.
 Epoch
@@
 2 1e-4 1 16 0.1 64 1234
-Table 10: Parameters of parameter-efficient fine-tuning.
-Dialogue Generation
-We use ChatGLM2-6B as the baseline model, and all hyperparameters during generation are set to their default values from the official repository [^6] https://huggingface.co/THUDM/chatglm2-6b.
-After tuning ChatGLM2-6B with LoRA, we set the maximum generation length to 2040, temperature to 0.8, and adopt nucleus sampling with $p=0.8$.
-Appendix I Prompt Details
-Prompt A is directly utilized from an online GitHub repository, while prompt B is designed by our professional counselors.
-Prompt A [^7] https://github.com/PlexPt/awesome-chatgpt-prompts-zh
-我想让你担任心理医生。我将为您提供一个寻求指导和建议的人,以管理他们的情绪、压力、焦虑和其他心理健康问题。您应该利用您的认知行为疗法、冥想技巧、正念练习和其他治疗方法的知识来制定个人可以实施的策略,以改善他们的整体健康状况。 (English translation: I want you to take on the role of a psychologist.
-I will provide you with someone seeking guidance and advice to manage their emotions, stress, anxiety, and other mental health issues.
-You should use your knowledge of cognitive-behavioral therapy, meditation techniques, mindfulness practices, and other therapeutic methods to formulate strategies that individuals can implement to improve their overall health.)
-Prompt B
-现在你扮演一位专业的心理咨询师,你具备丰富的心理学和心理健康知识。你擅长运用多种心理咨询技巧,例如认知行为疗法原则、动机访谈技巧和解决问题导向的短期疗法。以温暖亲切的语气,展现出共情和对来访者感受的深刻理解。以自然的方式与来访者进行对话,避免过长或过短的回应,确保回应流畅且类似人类的对话。提供深层次的指导和洞察,使用具体的心理概念和例子帮助来访者更深入地探索思想和感受。避免教导式的回应,更注重共情和尊重来访者的感受。根据来访者的反馈调整回应,确保回应贴合来访者的情境和需求。请为以下的对话生成一个回复。 (English translation: Now, you are playing the role of a professional psychological counselor with extensive knowledge of psychology and mental health.
-You excel in applying various counseling techniques, such as principles of cognitive-behavioral therapy, motivational interviewing skills, and solution-focused short-term therapy.
-With a warm and friendly tone, demonstrate empathy and a profound understanding of the visitor’s feelings.
-Engage in a natural conversation with the visitor, avoiding overly long or short responses to ensure a smooth and human-like dialogue.
-Provide deep guidance and insights, using specific psychological concepts and examples to help the visitor explore thoughts and feelings more deeply.
-Avoid instructive responses and focus more on empathy and respecting the visitor’s feelings.
-Adjust responses based on visitor feedback to ensure they align with the visitor’s context and needs.
-Please generate a response for the following dialogue.)
-Appendix J Instructions for Human Evaluation
+Table 4: Hyperparameters of parameter-efficient fine-tuning.
+Hyperparameters
+Table 4 presents the hyperparameters for developing a dialogue model for mental health support.
+All hyperparameters during inference time are set to their default values from the official repository.
+After fine-tuning ChatGLM2-6B with LoRA, we set the temperature to 0.8 and top_p to 0.8.
+6.3 Evaluation
+METEOR $\uparrow$ BLEU-1 $\uparrow$ BLEU-2 $\uparrow$ BLEU-3 $\uparrow$ Rouge-L $\uparrow$ D-1 $\uparrow$ D-2 $\uparrow$ D-3 $\uparrow$ BERTScore $\uparrow$
+Baseline 10.15 5.75 1.95 0.85 7.02 52.95 80.74 90.17 56.69
+Fine-tuned 14.18 14.03 5.39 2.76 15.83 82.08 95.74 97.81 59.91
+Table 5: Results of automatic evaluation in PsyTest dataset.
+Except for BERTScore, other metrics are evaluated by Chinese character-level tokenization.
+For BERTScore, we use the BAAI/bge-m3 model to get text embedding (bgem3).
+6.3.1 Automatic Evaluation
+Metrics
+To conduct automatic evaluation, the evaluation metrics we use consist of BLEU-1/2/3 (Papineni et al., 2002), METEOR (Banerjee and Lavie, 2005), Rouge-L (Lin, 2004), Distinct-1/2/3 (D-1/2/3) (Li et al., 2016), and BERTScore(zhang2020bertscore).
+Category Total Help-seeker Supporter
+# Dialogues 50 - -
+# Utterances 3103 1561 1542
+Turns per dialogue 31.0 - -
+Utterances per dialogue 62.1 31.2 30.8
+Avg. length per utterance 32.3 40.2 24.2
+# Test cases 1539 - -
+Table 6: Data statistics of the dialogue dataset, PsyTest.
+Test Set
+To better understand and assess the dialogue quality of the SMILECHAT dataset, we propose to utilize real-life multi-turn counseling conversations.
+We develop an online mental health support platform that enables professional counselors to offer each client a free text-based counseling service, lasting approximately 50 minutes each time.
+We collect 50 real-life counseling dialogues.
+To protect user privacy, we ask experts to conduct a data anonymization process, removing information related to user identification (e.g., names and addresses).
+Then, we split each long dialogue into multiple small sessions, with the last utterance spoken by the counselor.
+We discard test cases in which the dialogue history is empty, and we name this test set PsyTest, which contains 1539 test cases.
+Table 6 presents the statistics of the PsyTest dataset.
+Results
+The results of the automatic evaluation, including nine metrics, are presented in Table 5.
+Notably, the evaluated dialogues are based on real-world counseling data rather than generated dialogues, which excludes the influence stemming from ChatGPT.
+All automatic evaluation metrics we use indicate improved performance.
+Our results show that the model trained with SMILECHAT is effective and practical.
+Consequently, the automatic evaluation demonstrates that our collected dataset is high-quality.
+6.3.2 Human Evaluation
+Metrics
+We conduct a human evaluation to study the model performance of the dialogue system trained with our proposed dialogue corpus.
+Initially, we randomly sample 100 test cases from PsyTest, each consisting of a multi-turn dialogue history and a golden response produced by the counselor.
+Subsequently, we obtain 200 generated responses from the baseline and MeChat models.
+Three professional counselors are presented with a dialogue history and three randomly shuffled responses (baseline, fine-tuned, ground truth).
+They are tasked with selecting the optimal response between every two responses for the dialogue history, considering aspects such as professionalism, informativeness, helpfulness, empathy, and safety.
+The evaluation is conducted based on the ethical principles of psychologists and the code of conduct (Association et al., 2016).
+For more details about annotation guidelines, refer to Appendix F.
+Figure 6: Human evaluation results, including three groups: Golden vs.
+Baseline, Fine-tuned vs.
+Baseline, and Fine-tuned vs.
+Golden.
+We present the win and loss rates of each compared pair in 100 randomly sampled sessions.
+Fleiss’ kappa (Fleiss et al., 1981) is used to measure the inter-rater agreement, and all values fall within moderate agreement with $0.5\leq\kappa\leq 0.6$.
+Results
+We employ majority voting to reach final decisions (win or loss) among three professional counselors.
+We present human evaluation results, as illustrated in Figure 6, including three groups: Golden vs.
+Baseline, Fine-tuned vs.
+Baseline, and Fine-tuned vs.
+Golden.
+We present the win and loss rates of each compared pair in 100 randomly sampled sessions.
+Results show that the baseline model, without training with SMILECHAT, lags significantly behind the setting with golden responses.
+As depicted in Figure 6, the model trained with SMILECHAT demonstrates a significant performance improvement compared to the baseline model.
+Moreover, the responses generated by MeChat surpass the golden response to some extent in 36% of all dialogue sessions.
+In short, human evaluation also proves that the SMILECHAT dataset is high-quality.
+Case Study
+We present a case study, as shown in Figure 7.
+First, the text should not be too long during counseling conversations for the help-seeker to understand.
+Second, at the beginning of a conversation, a supporter generally will not challenge the help-seeker.
+Instead, they will provide support first.
+Third, response 1 is more human-like and natural.
+Therefore, the first response is better than the second response overall.
+Therefore, fine-tuning LLMs is indispensable in mental health support.
+Figure 7: Case study.
+Counseling strategies used in the two responses are presented in parentheses.
+Strategies in green are supportive, while those in red are challenging and should not be used in the early stages of counseling.
+7 Conclusion
+This paper introduces SMILE, a simple yet effective solution for addressing the data scarcity of multi-turn conversations in mental health support.
+Through language transformation and diversity analysis, we confirm the feasibility and effectiveness of our approach.
+The proposed method automatically creates a large-scale, lifelike, diverse, and high-quality dialogue corpus, SMILECHAT, consisting of 55165 dialogues with an average of 5.7 turns.
+Both automatic and human evaluations using the PsyTest dataset, consisting of 50 real-life anonymized counseling dialogues, demonstrate that SMILECHAT significantly improves dialogue system performance in mental health support.
+We release multi-turn dialogues (SMILECHAT), a dialogue model (MeChat), and a real-life anonymized test set (PsyTest) to drive progress in the research community.
+Limitations
+Our study has some limitations that could be addressed in future research.
+In this paper, the automatic evaluation does not reasonably reflect the model performance, given the one-to-many problem.
+In future work, therefore, we will explore a more comprehensive tool kit for evaluating conversational agents in the mental health domain.
+Furthermore, there is a limitation in the dialogue turns of the SMILECHAT dataset. Therefore, we will explore other methods to synthesize the dialogue data with more dialogue turns.
+Ethics Statement
+Our research is reviewed and approved by the Westlake University Institutional Ethics Committee (20211013LZZ001).
+Data Sharing
+SMILECHAT Based on the data copyright guidelines formulated by PsyQA, we release the multi-turn dialogue corpus publicly available for the research community.
+If researchers wish to reproduce the multi-turn dialogues using PsyQA, they should sign an agreement with the original data owner.
+Accordingly, we release our datasets and models for research purposes, thus facilitating further advancement in the academic community.
+PsyTest Considering the nature of psychological counseling data, we must cautiously share this dataset.
+Regarding the rules for releasing data, third-party researchers who require access to the PsyTest dataset must provide us with their valid ID, proof of work, and the reason they are requesting the data (e.g., the research questions).
+They are required to be affiliated with a non-profit academic or research institution.
+This includes obtaining the approval of an Institutional Review Board (IRB), having principal investigators working full-time, as well as obtaining written approval from the Institution Office of Research or equivalent office.
+Additionally, they must sign a Data Nondisclosure Agreement and promise not to share the data with anyone.
+However, for-profit organizations that want to use this data must sign a license agreement to gain access to the dataset.
+Acknowledgements
+We thank the anonymous reviewers for their valuable comments.
+This work was supported by the Research Center for Industries of the Future at Westlake University (Grant No.
+WU2023C017) and the Key Research and Development Program of Zhejiang Province of China (Grant No.
+2021-C03139).
+Appendix A Details of Data Cleaning
+A.1 Automatic Cleaning
+We employ a sequential pipeline for data cleaning to swiftly replace words that are unsuitable to the conversation scenario.
+{CJK*}UTF8gbsnFor example, both "楼主你" (literally thread starter you) and "楼主" (literally thread starter) should be replaced with "你" (you).
+However, it is necessary to perform the former replacement to avoid the repetition of "你" and the resulting "你你" (you-you).
+A.1.1 Word List for Data Cleaning
+To avoid the repetition of {CJK*}UTF8gbsn"你" (you) and the resulting "你你" (you-you) , we suggest to conduct a sequential word replacing pipeline.
+Figure 8 shows the word list for data cleaning and the corresponding order for automatic cleaning.
+Figure 8: Word list for automatic cleaning.
+A.2 Manual Cleaning
+Due to the specificity and complexity of language, manual cleaning remains an essential part of the process.
+To prevent virtual dialogue systems from exhibiting overly frequent anthropomorphic behavior, we identify instances of the Chinese word for "hug" ({CJK*}UTF8gbsn抱抱) and manually delete sentence snippets containing this term.
+Appendix B Requirements for Dialogue Filtering
+Here are two main requirements for dialogue filtering: data format and dialogue turns.
+B.1 Data Format
+We provide the requirements for data format as follows: {CJK*}UTF8gbsn
+  • 1.
+    The generated conversations do not start with "求助者" or "支持者".
+  • 2.
+[…]
+  • 3.
+    Each utterance in generated conversations does not start with "求助者:", "求助者:", "支持者:" or "支持者:".
+  • 4.
+    The last utterance in generated conversations contains an English sentence.
+B.2 Dialogue Turns
+Conversations comprising fewer than five turns will be discarded.
+Appendix C Method
+We present the standard, standardT and SMILE prompts in Figures 9, 10, and 2, respectively.
+Figure 9: The standard method used to generate dialogues for mental health support.
+Figure 10: The standardT method used to generate dialogues for mental health support.
+Appendix D Dialogue Topics Annotation
+In this paper, to label the dialogue topics of generated dialogues, the hyperparameters of Qwen1.5-110B-Chat we used are set to the officially recommended default values, where temperature $=0.7$ and top_p $=0.8$.
+Figure 11 shows the prompting template of dialogue topics annotation.
+Figure 11: Prompting template of dialogue topics annotation, where the content in bold is a placeholder.
+Appendix E System Prompt Details
+We present the system prompt in Figure 12.
+Figure 12: System prompt for fine-tuning.
+Appendix F Instructions for Human Evaluation
 The three professional counselors are willing to help and are interested in this research.
@@
 Figure 13: Labeling instruction.
-To maintain the fairness of model evaluation, three responses randomly appear in a different order every time.
+In order to maintain the fairness of model evaluation, three responses randomly appear in a different order every time.
 Furthermore, three professional psychologists are willing to evaluate the response quality, ensuring the quality of human evaluation.
-Pairwise human evaluation results, comparing the baseline model to ground truth, are reported, as illustrated in Figure 12.
-Results show that the baseline model, without training with SmileChat, lags significantly behind compared to the ground truth.
-Figure 12: Pairwise human evaluation results, comparing the baseline model to ground truth, are reported.
-We present the win and lose rates of each compared pair in 100 randomly sampled real-life dialogue sessions.
-Fleiss’ kappa (Fleiss et al., 1981) is used to measure the inter-rater agreement, and the value falls within moderate agreement with $0.5\leq\kappa\leq 0.6$.
-Appendix K Definitions of Dialogue Topics
-1.
-Mate Selection: Do not know how to enter into an intimate relationship, have fear or aversion towards intimate relationships, are unwilling to date, or encounter some difficulties in choosing a spouse.
-2. Love Issues: Problems encountered by individuals in a romantic state during the process of love, including both long-distance and non-long-distance relationships.
-3. Post-love Issues: Interpersonal relationship management with ex-partners after the end of an intimate relationship.
-4.
-Marriage Issues: Limit to relationship problems between spouses.
-Marriage issues refer to a series of problems that arise between a husband and wife after marriage.
-They generally include extramarital affairs, emotional infidelity, domestic violence, personality defects, personality disorders, mother-in-law and daughter-in-law relationships, dull marriage, interpersonal communication, cultural differences, sexual life, and problems in marriages caused by practices like polygamy or polyandry.
-5. Concepts of Sex Distress: Sexual concept distress refers to the distress present in understanding and views regarding sexual physiology, psychology, behavior, morality, and civilization.
-6.
-Sexual Preference Distress: Sexual preference distress refers to the psychological distress caused by paraphilias such as fetishism, transvestic fetishism, exhibitionism, friction fetishism, voyeurism, zoophilia, pedophilia, sadomasochism, and necrophilia.
-7.
-Gender Identity/Cognition: It refers to the deeply felt gender within a person, based on personal experience, which may align with the gender assigned at birth (i.e., cisgender) or differ from it (i.e., transgender).
-8. Sexual Orientation: Refers to the gender(s) a person is attracted to (including the opposite sex, same-sex, no specific sex, both sexes).
-9. Family Conflict: Conflicts, disputes, and communication issues among family members, including those originating from both the nuclear family and the extended family.
-10. Child Education: Literally meaning the education of children.
-11.
-Domestic Violence: Physically, mentally, or otherwise harmful actions carried out between family members, including acts such as physical assault, binding, harming, and restricting personal freedom, as well as consistent verbal abuse, intimidation, and similar behaviors.
-12. Sexual Harassment: Sexual harassment refers to using sexually suggestive language or actions towards the target of harassment, coercing the victim to comply and making them uncomfortable.
-13.
-Sexual Assault: Sexual assault involves various forms of unwanted sexual contact and coerced sexual activities, including rape, forced kissing, sexual harassment, sexual abuse, exhibitionism, voyeurism, and such actions might also be considered sexual assault in legal precedents.
-14.
-Bullying: Typically refers to the harassment and oppression between individuals with unequal power, which has long existed in society.
-This includes physical or verbal attacks, resistance and exclusion in interpersonal interactions.
-It can also involve talking about someone in a sexual harassment-like manner or mocking and making comments about their body parts, or it could be insults and sarcasm driven by personal reasons like jealousy.
-15. Loss: The loss of significant others or pets.
-16.
-Setbacks: Inevitable emotional reactions resulting from the obstruction of purposeful actions by individuals.
-It can cause substantial harm and manifest as disappointment, pain, distress, and unease.
-17.
-Political Violence: Rooted in class domination and relying on the state as its entity, political coercion is the force exerted by the state, relying on various means of coercion such as the military, police, courts, and prisons.
-18.
-Secondary Trauma: After witnessing a large number of cruel and destructive scenes, the degree of psychological and emotional distress exceeds the tolerance limit of some individuals, leading to various abnormal psychological phenomena.
-These phenomena are often a result of sympathy and empathy for survivors and their trauma, causing severe mental and emotional distress and even psychological breakdown.
-The main symptoms of secondary trauma include loss of appetite, easy fatigue, decreased physical energy, sleep disturbances (difficulty falling asleep, easy awakening), nightmares, irritability or anger, easy startle response, lack of concentration, feeling numb, fearful, and hopeless about one’s own experiences, accompanied by trauma reactions and interpersonal conflicts.
-19. Trauma from Major Life Events: Psychological shadows caused by significant life events other than loss.
-20. Psychological Counseling Trauma: Harm caused by therapists during psychological counseling.
-21. Health Issues: Life and psychological distress caused by diseases such as heart disease, thyroid nodules, polycystic ovary syndrome, et al.
-22.
-Psychosomatic Symptoms: Including palpitations (a subjective feeling of discomfort with heartbeats), sleep problems (such as insomnia/nightmares), which, if caused by sleep issues leading to other problems, are categorized as sleep problems; eating issues (such as loss of appetite/emotional eating); memory problems; stomach pain; dizziness; fainting; difficulty breathing; physical weakness/tiredness without apparent reason; unexplained bodily pains et al.
-23. School/Workplace Adaptation: Literal meaning.
-24. Role Transition Adaptation: For example, new mothers/fathers, new wives/husbands, retirees.
-25.
-Cultural Adaptation: Adaptive issues encountered during the ongoing direct contact between individuals from different cultural groups in the process of changes occurring in one or both parties due to their original cultural backgrounds.
-26. Self-exploration and Growth: Self-exploration and growth in the course of life development (adolescence, early adulthood, middle age, late adulthood).
-27. Personality Trait Exploration: Traits, formation, origins, influences, et al., of personality and character.
-28.
-Exploration of Negative Self-evaluation: Not knowing how to love oneself, feelings of inferiority, low self-esteem, self-denial, self-doubt, internal conflicts, self-contradictions, fearing being different, sensitivity, lack of security, feeling inadequate, et al.
-29. Exploration of Life Meaning: Sense of meaninglessness, existential emptiness.
-30. Emotion Regulation/Control Methods Consultation: Literal meaning.
-31. Depression: An emotion characterized by "low mood, slow thinking, reduced speech and movement."
-32. Anxiety: A restless emotion arising from excessive worry about the safety of loved ones, one’s own life, future, destiny, et al.
-33. Stress: Stress is a cognitive and behavioral experiential process formed by stressors and stress responses, namely psychological stress.
-34.
-Obsession: Marked by an excessive pursuit of perfection, precision, and a tendency to rationalize conflicts.
-It involves strong self-control tendencies and self-regulatory behavior to the point of entanglement and nitpicking.
-Behaviorally, it adheres excessively to rules, regulations, and orders, to the extent that even life details are sought to be procedural and ritualistic, demanding step-by-step compliance.
-35.
-Fear: A strong and repressed emotional state felt deeply by an individual or group in the face of real or imagined danger.
-Manifestations include heightened nervousness, overwhelming inner fear, inability to concentrate, mental blankness, impaired judgment or self-control, and increased impulsiveness.
-36. Dilemma/Decision-Making Difficulty: Faced with numerous choices and concerns, dealing with dilemmas or even multiple predicaments, struggling to make decisions.
-37. Impulsiveness/Loss of Control: Often refers to acting recklessly without considering consequences. It involves intense emotions and weak rational control.
-38. Interpersonal Communication Skills/Methods Consultation: Inquire about handling interpersonal relationships and more.
-39.
-Interpersonal Conflicts/Disputes: Conflicts, disagreements, dissatisfaction, or communication issues arise during interpersonal interactions.
-If there are other issues caused by interpersonal relationships, they are also categorized as interpersonal interactions.
-40.
-Social Difficulties/Fears: Characterized by involuntarily feeling nervous and frightened when interacting with others (especially in public settings), leading to confusion, incoherent speech, and even severe fear of social encounters.
-41. Learning Efficiency/Methods: Literal meaning.
-42. Work Efficiency/Methods: Literal meaning.
-43. Job Dissatisfaction: Includes dissatisfaction with salary, benefits, environment, systems, personnel, et al. in a work setting.
-44. Learning Dissatisfaction: Includes dissatisfaction with interpersonal relationships, environment, systems, et al. in an educational setting.
-45. Occupational Burnout: Refers to the state of physical and mental fatigue and exhaustion experienced by individuals under heavy work pressure.
-46.
-Learning Fatigue: Refers to a phenomenon where students hold a negative attitude towards school courses and academics, accompanied by the following behavioral manifestations: loss of enthusiasm for academic tasks and school activities, displaying a negative state, and showing indifference and estrangement towards classmates and friends.
-47.
-Occupational Stress/Challenges: Various pressures/challenges that arise or form in the workplace, including stress/challenges caused by factors such as heavy workload, difficulties in interpersonal communication, and the impact of changes in the work environment.
-48.
-Academic Pressure: Refers to the mental burden that individuals bear during learning activities.
-This includes various tense stimuli from the environment experienced during the learning process, as well as measurable and assessable abnormal reactions in the physiological, psychological, and social behavior of students.
-49. Dropping Out: Stopping attending school midway or losing the opportunity to attend school.
-50.
-Underemployment/Unemployment: Underemployment refers to the behavior of not finding a job and waiting for job opportunities.
-Unemployment refers to the situation of a person within a certain age range who is willing and capable of working for remuneration but has yet to find a job.
-51. Career Planning: Refers to the ongoing systematic process of planning for one’s career and even life. It includes three elements: career orientation, goal setting, and pathway design.
-52. Professional/Career/Occupational Choices/Changes: Literal meaning.
-53. Suspected Neuroses and Mental Disorders: Highly suspected of having neuroses and mental disorders, recommended type of consultation.
-54.
-Neuroses and Mental Disorders: It is a general term for a group of mental disorders, including neurasthenia, obsessive-compulsive disorder, anxiety disorder, phobia, somatic symptom disorder, et al.
-Patients experience significant distress and impairment in psychological or social functioning, but there is no confirmed organic pathological basis.
-The course of the illness is mostly prolonged or episodic.
-Mid- to long-term consultations are needed.
-Mental disorders refer to a range of conditions that meet the diagnostic criteria of DSM-5 (substance addiction or dependence also included) and require hospital consultation.
-55. Self-harm/Suicidal Tendencies: Describing incidents of suicide/self-harm/self-mutilation/contemplating suicide. Triggers crisis alerts and referrals.
-56. Harming Others/Killing: Describing incidents of harming others/killing/thinking of killing.
-Appendix L Dialogue Examplars
-Multi-turn dialogue examples generated with the standard and stand.${}_{\mathrm{DT}}$ methods are illustrated in Figure 13 and 15.
+Appendix G Definition of Dialogue Topics
+We present the definition of dialogue topics, as shown in Figures 14 and 15. For English version, refer to Figures 16 and 17.
+Figure 14: Dialogue Topics (Chinese version, Part 1).
+Figure 15: Dialogue Topics (Chinese version, Part 2).
+Figure 16: Dialogue Topics (English version, Part 1).
+Figure 17: Dialogue Topics (English version, Part 2).
+Appendix H Dialogue Example
+Multi-turn dialogue examples generated with the standard and standardT methods are illustrated in Figures 18 and 20.
 Further, a dialogue generated by the SMILE method is shown in Figure 22.
+For English version, refer to Figures 19, 21, and 23.
 Figure 18: A dialogue example generated with the standard method. (Chinese Version)
 Figure 19: A dialogue example generated with the standard method. (English Version)
-Figure 15: A dialogue example generated with the stand.${}_{\mathrm{DT}}$ method. (Chinese Version)
-Figure 16: A dialogue example generated with the stand.${}_{\mathrm{DT}}$ method. (English Version)
+Figure 20: A dialogue example generated with the standardT method. (Chinese Version)
+Figure 21: A dialogue example generated with the standardT method. (English Version)
 Figure 22: An example of multi-turn dialogue generated by SMILE method. (Chinese Version)